### v1

# Secure Human Oversight of AI: Exploring the Attack Surface of Human Oversight

Abstract
Human oversight of AI is promoted as a safeguard against risks such as inaccurate outputs, system malfunctions, or violations of fundamental rights, and is mandated in regulation like the European AI Act. Yet debates on human oversight have largely focused on its effectiveness, while overlooking a critical dimension: the security of human oversight. We argue that human oversight creates a new attack surface within the safety, security, and accountability architecture of AI operations. Drawing on cybersecurity perspectives, we analyze attack vectors that threaten the requirements of effective human oversight, thereby undermining the safety of AI operations. Such attacks may target the AI system, its communication with oversight personnel, or the personnel themselves. We then outline hardening strategies to mitigate these risks. Our contributions are: (1) introducing a security perspective on human oversight, and (2) providing an overview of attack vectors and hardening strategies to enable secure human oversight of AI.
Keywords Human oversight  $\cdot$ cybersecurity  $\cdot$ attack vectors  $\cdot$ hardening strategies  $\cdot$ AI governance
1 Introduction
Figure 1: The human oversight attack surface encompasses the AI system’s technical infrastructure, the human oversight personnel, and the communication channels between the two sides. Attacks on human oversight aim to undermine oversight personnel’s epistemic access, causal power, self-control, and fitting intentions, the four requirements of effective human oversight [1]. Malicious actors can utilize different attack vectors to undermine one or several of these requirements (see Table 1). In order to protect human oversight from malicious actors, various hardening strategies need to be considered and deployed to counteract the threats posed by the attack vectors (see Table 2). A well-designed combination of hardening strategies can shield against different attack vectors and, thus assure secure human oversight.
Human oversight of artificial intelligence (AI) is often considered a strategy to address risks associated with the use of AI in high-risk contexts [2, 3, 4]. Human oversight personnel are expected to detect inaccurate outputs, outputs that may pose risks to fundamental human rights (e.g., discriminatory outputs), unsafe states of AI systems, and system malfunctions, and to intervene in order to mitigate or minimize these risks [5]. Human oversight is also a key concept in AI governance and ethics [3, 2] and a central requirement for the implementation of high-risk AI systems in emerging AI regulation, such as Article 14 of the European AI Act [6, 3]. However, the idea of relying on humans to oversee AI in order to mitigate operational risks also faces criticism and is expected to encounter substantial challenges in practice [7, 2, 1]. The most prominent criticisms and challenges include doubts about whether humans can effectively oversee AI systems [2], the difficulty of defining what constitutes effective human oversight [1], clarifying who should provide oversight and when it should occur [6], specifying which risks and unwanted system behaviors oversight personnel should detect [5], and determining how to design and evaluate effective human oversight [1].
However, one key challenge that remains underexplored is the security risks that arise from human oversight of AI. Because human oversight will be implemented as part of the safety, security, and accountability architecture of AI systems [3, 7], malicious actors who successfully attack elements related to oversight can significantly undermine the safety of AI operations. In this sense, the human oversight architecture creates a new attack surface for targeting AI operations in high-risk contexts such as critical infrastructure, public administration, or medicine. A security-focused perspective on human oversight is crucial, given the central role oversight already plays, the expectations placed on it in high-risk contexts, and its growing importance as a regulatory requirement in AI governance, such as under the EU AI Act. If we fail to assess and address the security challenges of human oversight, we risk overlooking potential attack vectors in high-risk contexts; and without awareness of these vectors, we will not be able to design effective hardening strategies.
In this paper, we introduce a security-related perspective on human oversight of AI. Specifically, drawing on cybersecurity research, we outline possible attack vectors that threaten the key requirements of effective human oversight (i.e., epistemic access, self-control, causal power, and fitting intentions; [1]). These attacks may target the AI system itself, the communication between the AI system and oversight personnel, or the personnel directly (see Figure 1). We also draw on the cybersecurity literature to discuss hardening strategies that can mitigate the risks posed by such attacks.
This paper makes two key contributions. First, we introduce a security-related perspective on human oversight of AI. This perspective is critically missing from current research, even though human oversight is becoming an integral component of the safety, security, and accountability architecture of AI operations, and thus an increasingly viable target for malicious actors. Second, we provide an overview of possible attack vectors and hardening strategies contextualized for human oversight of AI. We hope this contribution supports AI governance and system design in ensuring not only effective but also secure human oversight of AI.
2 Related Work
2.1 Effective Human Oversight of AI
Research discusses the nature of human oversight of AI, including the what, when, and who of oversight. Regarding the what, many forms of human-AI collaboration in principle involve some degree of human oversight. Across a wide range of human-AI interaction use cases, humans monitor AI system outputs, check their accuracy, request explanations, adapt outputs, intervene when systems malfunction, or take over manual control in the event of failures [8, 9, 10, 11, 12, 13, 14, 15]. In a more targeted sense, and following how emerging research has tentatively defined the concept [1], human oversight specifically refers to humans mitigating risks in AI operations [16, 6]. More concretely, oversight personnel are tasked with monitoring AI systems, detecting failures and malfunctions, identifying unsafe states, flagging inaccurate outputs, and recognizing outputs that may pose risks to fundamental human rights, such as discriminatory outputs [1, 5, 6]. Once such issues are detected, oversight personnel are expected to act in ways that effectively mitigate emerging risks [5, 1]. This can involve taking over manual control, overwriting inaccurate or inadequate outputs [17], or stopping AI operations and triggering intervention by other human stakeholders, such as developers who can adapt the system. Importantly, the AI system under oversight may be operated by another human who is not considered oversight personnel but a user of the system. For example, a patient using an AI chatbot for mental health support [18] would be the user, while oversight personnel monitor the patient–AI interaction to detect risky situations (e.g., if the patient discusses suicidal behavior with the chatbot).
Regarding the when of human oversight, human involvement is important at every stage of AI development: during system design, at run time, and when inspecting errors and failures in hindsight [4, 19, 20]. In a more targeted sense, human oversight usually refers to personnel overseeing AI systems during run time [1, 3]. This can involve dynamic and time-critical situations, such as humans overseeing autonomous vehicles [17, 21], but also less dynamic contexts, such as AI in judicial decision-making or hiring, where there is more time to deliberate about the accuracy or adequacy of AI outputs [5, 22].
Regarding the who of human oversight, any human with the skills necessary to effectively oversee an AI system could serve as oversight personnel [6, 23]. These individuals may be domain experts such as doctors or judges, who can evaluate the adequacy of specific AI outputs based on their expertise. Alternatively, oversight personnel may be trained to detect AI failures without being domain experts in the task performed by the AI. For example, in a medical AI setting, one oversight person might monitor the system’s technical functions on an aggregate level, while medical professionals evaluate individual outputs requiring domain knowledge [23, 5]. Human oversight can also be performed by a single individual or by a team collaboratively overseeing AI systems.
Research has also examined the requirements for effective human oversight. Sterz et al. [1] identify four key aspects: epistemic access, causal power, self-control, and fitting intentions. As we discuss possible attack vectors, we consider how successful attacks might undermine these requirements. Epistemic access means that oversight personnel possess sufficient knowledge about the AI system and understand the consequences of their interventions. Without it, they may fail to detect system failures or inaccurate outputs. Causal power refers to the ability to intervene in ways that actually influence AI processes and outputs. Without causal power, oversight personnel may be unable to overwrite or correct system outputs. Self-control requires oversight personnel to maintain agency over their own actions. For instance, fatigue or intoxication may impair their capacity to fulfill oversight duties. Likewise, being threatened by another person can also undermine self-control, as the actions performed by the oversight personnel are then steered by another agent. Fitting intentions means that personnel are motivated to carry out their responsibilities with the right intentions. In an extreme case, a malicious actor might be hired as oversight personnel and deliberately undermine AI safety. In a less extreme case, poor working conditions may erode motivation and attention, weakening oversight effectiveness [1].
Finally, regarding how to design effective human oversight, insights from the broader field of human-AI collaboration provide valuable guidance (e.g., [8, 24, 16]). Research emphasizes the sociotechnical nature of oversight and highlights the need to design technical, human, and contextual factors that support effective oversight [5, 3, 1]. Technical factors include system interface design, explainability approaches, and interaction modes between humans and AI systems [9, 25, 26, 27]. Human factors include the skills oversight personnel need and the training they receive [23, 16, 5]. Contextual factors include the design of the working environment, time pressure, and the clarity of roles and responsibilities within oversight teams [5, 28, 29].
2.2 Attack Vectors and Hardening Strategies in Cybersecurity
With the implementation of AI systems across diverse fields of application, ensuring the security of deployed solutions has become increasingly urgent. Achieving this requires an in-depth understanding of specific attack vectors and the attack surface within each application domain [30]. Traditionally, an attack vector is defined as a method or pathway that a malicious actor uses to gain access to a network or system. In this paper, we extend the definition to also include methods or pathways that undermine the effectiveness of human oversight. We use the term attack surface to describe the total set of points where a malicious actor could attempt to enter, exploit, or otherwise attack digital and AI infrastructures [30]. For example, research has systematically analyzed attack vectors in autonomous driving systems and defined the corresponding attack surface [31]. Other work has examined the attack surface in federated learning workflows [32]. For human oversight, however, such systematic investigations of the attack surface are still missing.
Relevant related work can be found in the field of AI cybersecurity, which examines attack vectors that explicitly target AI systems. For instance, poisoning attacks manipulate training data to introduce backdoors into AI models that can later be exploited [33, 34, 35]. Adversarial attacks embed imperceptible patterns into inputs to change model behavior, often causing misclassifications or other problematic outcomes [36, 37, 38]. Another line of research targets explainable AI (XAI) methods designed to enhance understanding of AI decisions [25]. Such attacks manipulate explanations to obscure the real reasons behind a model’s decision [39, 40, 41].
Because human oversight extends beyond an isolated AI system, additional relevant research comes from traditional cybersecurity. Distributed denial-of-service (DDoS) attacks flood networks with traffic, making them unresponsive or unavailable [42, 43]. Man-in-the-middle attacks intercept and manipulate data traffic between two endpoints [44]. Another common vector is malware, which – due to its diversity – can be deployed against nearly any digital system [45, 46].
Particularly relevant for defining the attack surface are attacks on the human factor in cybersecurity. Social engineering exploits human vulnerabilities to extract sensitive information or gain access privileges [47, 48]. Two additional but in cybersecurity research less discussed vectors are coercion and bribery, which can likewise be used to gain sensitive information or access privileges [49, 50]. Finally, insider threats exploit individuals inside an organization (e.g., disgruntled employees) to carry out attacks [51]. With the growing threat of cyberattacks, every system that incorporates digital components must be analyzed for potential attack vectors. Our work contributes to this literature by contextualizing attack vectors for the specifics of human oversight of AI.
In response to the many ways digital systems can be attacked, cybersecurity research has developed hardening strategies as methods designed to secure systems against attacks [52, 53, 54]. One line of work establishes methods for securing data traffic in computer networks. Network management systems can detect malicious traffic [55, 56], while encryption protects legitimate communications [57, 58]. To monitor organizational IT systems, intrusion detection systems have been introduced and are continuously improved [59, 60]. Another widely used practice is red teaming, where every system component is rigorously tested by simulating attacks; research continues to optimize red teaming approaches [61, 62]. Beyond technical strategies, organizational measures also play a key role. Transparency of software components is crucial for analyzing systems and identifying potentially harmful functionalities [63, 64]. Other well-established approaches include governance principles and audit frameworks that help secure deployed systems [65, 66]. Finally, like attack vectors, hardening strategies must also address the human factor. Research has explored how to improve system security by training users and raising awareness of risks [67, 68].
In this paper, we extend the literature by contextualizing hardening strategies specifically for securing human oversight of AI against malicious attacks.
3 The Human Oversight Attack Surface
Human oversight can serve as part of the safety, security, and accountability architecture of AI systems, as oversight personnel are expected to detect and mitigate risks. In addition, oversight personnel may be required to actively confirm that an AI system is in a safe state. Human oversight may become even more important in AI operations if it evolves into a central requirement of AI governance and regulation in high-risk contexts [2, 6, 3, 1].
With a more wide-spread implementation, human oversight will also become an increasingly attractive target for malicious actors seeking to corrupt or disrupt AI-based operations. Such actors could be cybercriminals targeting high-risk AI systems for ransom or foreign governments attacking a country’s critical infrastructure in hybrid or open conflict. Furthermore, sufficiently powerful AI agents that are successfully attacked could themselves become malicious actors by attempting to evade oversight [69, 70]. Considering these threats, designing human oversight requires not only ensuring effective oversight [1] but also ensuring secure oversight.
Designing for secure human oversight demands an understanding of its attack surface. The following list of attack vectors inspired by the broader cybersecurity literature is not exhaustive but illustrates the breadth of the human oversight attack surface. It outlines attacks that target the four core requirements of effective human oversight: epistemic access, self-control, causal power, and fitting intentions. Table 1 summarizes these attack vectors and their relation to the requirements for effective human oversight.
Table 1: Attack vectors that target the effectiveness of human oversight. For each attack vector, we indicate the specific requirement for effective human oversight that the attack seeks to undermine.
Attack vectors Affected requirements of effective human oversight
Poisoning attacks Epistemic access, self-control
Adversarial attacks Epistemic access
Explainability attacks Epistemic access
(Distributed) denial of service attacks Causal power
Man-in-the-middle attacks Epistemic access, causal power
Vulnerability exploitation Epistemic access, causal power
Malware Epistemic access, causal power
Social engineering Epistemic access, self-control
Coercion Self-control
Bribery Fitting intentions
Insider threats Epistemic access, causal power, fitting intentions
3.1 Poisoning Attacks
Poisoning attacks target the core functionality of AI models, namely the association of patterns in data with outcomes (e.g., real-valued numbers in regression, class probabilities in classification, or new text/image data in generative AI) [71]. These attacks exploit the fact that modern AI models are often trained on large volumes of (semi-)automatically collected data, for example through web crawlers. Malicious actors can insert triggers, or predefined patterns, into such datasets. When models are trained on compromised data, they may consistently exhibit malicious behaviors whenever a trigger is presented [72].
Poisoning attacks can compromise both the epistemic access and the self-control of oversight personnel. They undermine epistemic access because poisoned samples containing hidden triggers cause AI responses that deviate from normal outputs. A carefully crafted trigger, such as an imperceptible pixel pattern in images [73], can prevent oversight personnel from recognizing the behavioral shift. Even when oversight personnel detect that something is amiss, hidden triggers may still make it difficult to understand the malicious behavior.
Poisoning attacks can also erode self-control. If triggered outputs cause the oversight interface to generate excessive micro-notifications or confirmation requests, malicious actors can exploit this to induce fatigue by overwhelming oversight personnel with trivial but time- and energy-consuming tasks. This not only reduces oversight efficiency by consuming resources but also increases the likelihood of errors once fatigue reaches a critical level, thereby undermining the security of human–AI interaction.
3.2 Adversarial Attacks
Adversarial attacks aim to induce misclassifications in the outputs of an AI model [71]. By feeding the system with adversarial examples, attackers seek either to fool the model or to reduce its accuracy and performance. These adversarial examples typically resemble normal inputs but include subtle perturbations that may not be perceptible to oversight personnel, leading to misclassifications aligned with the attacker’s goals [72].
Adversarial attacks primarily compromise the epistemic access of oversight personnel. A crucial part of understanding AI behavior is knowing how input patterns are mapped to outputs (e.g., which structures or features in an input image correspond to specific output classes). Without this knowledge, reliably determining whether the system behaves acceptably becomes exceedingly difficult. Adversarial perturbations obfuscate the relationship between inputs and outputs, reducing epistemic access to the model’s true behavior. As a result, oversight personnel may fail to detect unwanted model behavior or even system failures.
3.3 Explainability Attacks
Attacks on explainability specifically target explainability and interpretability approaches and tools. These attacks manipulate or deceive the evidence for a model’s reasoning. A key characteristic of such attacks is their ability to alter the explanation while keeping the model’s output unchanged. This means that the AI system may still produce the same output, but the explanation for that output is inaccurate or misleading [74].
Explainability attacks primarily undermine the epistemic access of oversight personnel. Misleading or inaccurate explanations prevent personnel from understanding the AI system’s true behavior. Attackers can exploit these attacks to fabricate behavior that serves their goals. Even with extensive querying, oversight personnel may be unable to detect such manipulations [75]. Moreover, a mismatch between outputs and explanations adversely affects personnel who rely on explanations to identify problematic AI behavior [25, 76].
These attacks can also induce a state of deceptive epistemic access, where the system appears transparent but provides inaccurate explanations. This scenario can be more dangerous than a “black box” system because it fosters unjustified trust and makes detecting incorrect outputs more difficult. Consequently, a key focus of explainability for human oversight must be ensuring the fidelity and robustness of explanations [77, 78].
3.4 (Distributed) Denial of Service Attacks
Denial of service (DoS) attacks refer to methods that render a system inaccessible [79]. The core principle of DoS attacks is to overwhelm a target system with more requests than it can handle, causing the system to become nonfunctional. Various techniques exist to perform a DoS attack, such as exploiting the TCP/IP protocol (SYN flooding), using the ping functionality (ping flooding), or targeting mail servers (mail flooding). A variation of DoS attacks is the distributed denial of service (DDoS) attack, which distributes requests across multiple systems rather than originating from a single source [80].
(Distributed) DoS attacks can undermine the causal power of oversight personnel. Modern computing systems are increasingly implemented as network-based, distributed systems—a standard likely to apply in human oversight settings, where the AI system runs on a cloud-based or on-premise server while oversight personnel interact with it via network protocols. (D)DoS attacks can make either the AI system or the human oversight interface inaccessible. This threat is especially severe for cloud-based AI systems: because the AI system and oversight personnel are physically separated, malicious actors can effectively disable oversight through (D)DoS attacks while the AI system continues to operate in an unsafe state or produces inaccurate outputs.
3.5 Man-in-the-Middle Attacks
An attack vector that can be used in computer networks is the Man-in-the-Middle (MitM) attack [44]. When successful, this strategy allows malicious actors to intercept and control data traffic between two or more systems, granting them access to the exchanged information and enabling them to manipulate it. Additionally, attackers can conceal their presence, tricking the involved systems into believing they are communicating securely with each other [44]. Various methods exist to carry out a MitM attack: ARP spoofing can be used when the attacker and target are on the same local area network (LAN), snarfing can be applied in WLAN-based attacks, and domain name system (DNS) cache poisoning is another possible approach.
MitM attacks can compromise the epistemic access and causal power of oversight personnel, depending on which direction the manipulated information travels. If the attacker intercepts and alters information from the AI system to the oversight personnel, epistemic access is undermined: the true behavior, functionality, or outputs of the AI system may remain hidden.
If the manipulated information flows from the human to the AI system, causal power is compromised. For example, if an attacker modifies commands issued by oversight personnel, they lose the ability to interact effectively with the AI system and cannot intervene in AI processes. This situation is particularly dangerous because the attacker can send manipulated responses back to the personnel, creating an illusion of control and misleading them into believing they are still effectively supervising the AI system.
3.6 Vulnerability Exploitation
Vulnerability exploitation refers to utilizing existing flaws in hardware and software for attacks, such as remote code execution or exfiltration of sensitive data. Since modern digital systems are complex combinations of hardware and software components and their respective supply chains, flaws can accumulate quickly and pose a significant threat. One of the most dangerous vulnerability exploitation is the zero-day exploit [81]. This term refers to the exploitation of a specific vulnerability within a hardware or software component that is unknown even to the manufacturer of the component. This gives the attacker the time advantage of day zero, before the manufacturer can detect, identify and fix the vulnerability.
Attacks based on vulnerability exploitation can target epistemic access or causal power. Attackers can use the vulnerabilities to get unauthorized access to the system and in this way manipulate the system output which the human oversight person uses to detect deviations in system behavior. If the attacker can obfuscate their actions from the human oversight personnel, the personnel are prevented from understanding the real functioning of the AI system. Similar effects can be achieved by manipulating other information that human oversight personnel uses to assess the AI system.
Causal power is affected by vulnerability exploitation if attackers prevent human oversight personnel from appropriately communicating with the AI system. This can be achieved by dropping commands send by the human oversight personnel or disabling crucial components of the human oversight software. Depending on the exploitation, attackers might be able to disguise the attack. For example, and similar to the example for MitM attacks, they could drop commands by human oversight personnel but fake a response by the AI system that suggests the command was successfully executed.
3.7 Malware
Malware is software that executes harmful functionality on an IT system, often without the user’s knowledge. Its potential targets and effects are highly diverse. Ransomware encrypts the file system and typically demands a ransom for decryption. Spyware extracts information from the infected system. These are just two of the most common types, but malware encompasses a wide range of threats designed for various purposes [82, 45].
Due to the diversity of malware functionality, it can compromise epistemic access and causal power of oversight personnel. Regarding epistemic access, malware can alter the core functionality of an AI system. For example, it could insert system prompts into deployed large language models (LLMs), significantly changing their behavior [83]. Malware can also intercept or manipulate data traffic from the AI system to oversight personnel, or modify, encrypt, or obfuscate system logs. These actions prevent personnel from understanding the system’s true functionality and therefore undermine epistemic access.
Causal power can be compromised if malware intercepts or manipulates data traveling from oversight personnel to the AI system. For instance, if malware prevents personnel from interacting with the AI system or locks them out of controlling its core functions, they lose the ability to intervene in system processes or outputs. Similarly, if malware secretly alters commands issued by oversight personnel, it further undermines their causal power.
3.8 Social Engineering
Social engineering refers to attacks that exploit human characteristics [84]. These attacks leverage traits such as fear, respect for authority, or the willingness to help, tricking victims into revealing personal information, disabling security measures, or installing malware. The widespread use of digital communication channels has provided malicious actors with effective and scalable tools for conducting social engineering attacks [85].
Social engineering can compromise epistemic access and self-control of oversight personnel. If an attacker impersonates individuals with technical expertise or authority (e.g., trusted colleagues, management, external experts) they can distort personnel’s understanding of relevant information needed to oversee AI systems. For example, an attacker might pretend to be a technician from the AI provider, convincing human oversight personnel that certain graphical user interface notifications indicate a bug. This could lead personnel to ignore important alerts or install malware disguised as a hotfix. Similarly, attackers impersonating management may convince oversight personnel that specific system notifications are non-critical, framing them as part of corporate processes. In both cases, social engineering feeds false information about the AI system’s functionality to oversight personnel, thereby undermining epistemic access.
Social engineering can also erode self-control. By impersonating senior colleagues or management, attackers can exploit perceived authority to pressure oversight personnel into taking specific actions. If the attack is credible, personnel might follow instructions against their better judgment, believing the directives come from a legitimate authority.
3.9 Coercion
Coercion is a broad concept that extends beyond IT security and involves using force, often through threats or violations, to influence a party. In the context of human oversight, coercion refers to pressure or force that alters the intended actions of oversight personnel.
Coercion undermines the self-control of oversight personnel. By influencing oversight personnel to act according to the instructions of a malicious actor rather than their own judgment, coercion renders human oversight ineffective. Even when oversight personnel detect a need for intervention, they may be unable to act according to their assessment. In extreme cases, a malicious actor could threaten the oversight personnel to gain control over their actions, effectively nullifying the intended benefits of human oversight.
3.10 Bribery
Bribery refers to the act of persuading a person, organization, or authority to act in accordance with the goals of a malicious actor, typically by offering monetary incentives, status, or similar benefits.
Bribery undermines the fitting intentions of human oversight personnel. Usually, the intentions of human oversight personnel should align with fulfilling their oversight duties, such as detecting inaccurate outputs or system failures [1]. A successful bribe can misalign these intentions, causing personnel to act in accordance with the malicious actor’s goals instead. Such goals might include damaging critical infrastructure or compromising the functioning of AI in high-risk applications, such as in hospitals.
3.11 Insider Threats
Insider threats encompass attack scenarios in which malicious actors infiltrate the organization they intend to compromise [86]. Such attacks share the common characteristic that the attacker either has direct access themselves or exploits an employee’s access rights, whether voluntarily or accidentally. Two pathways to insider threats are discussed under the sections on coercion (see 3.9) and bribery (see 3.10). Another common pathway is the exploitation of disgruntled employees.
Insider threats can undermine epistemic access, causal power, and fitting intentions of oversight personnel. They can erode epistemic access by altering or falsifying monitoring data from the AI system. Since human oversight personnel rely on this data to accurately assess AI behavior and intervene when necessary, this prevents them from understanding the system. Insider threats can also obfuscate AI system documentation by deleting or modifying key information, further impeding epistemic access.
Insider threats can also compromise the causal power of oversight personnel. For example, possible sequences of actions that can be taken as a response to detecting problematic system behavior can be ambiguous and the effectiveness of different sequences of actions can differ vastly. Attackers may be able to change organizational oversight policies on how to deal with inaccurate outputs or system malfunctions. In the worst case, such adapted policies can prevent oversight personnel from acting appropriately after detecting problematic system behavior which negatively affects causal power.
Another example of how insider threats may undermine causal power is by achieving and thereby exploit under-staffing. Redirecting resource allocation away from human oversight or preventing human oversight personnel from doing their work (e.g., by force) can reduce the workforce available to human oversight. While this requires substantial access and intervention capabilities, it can prevent oversight personnel from exercising causal power effectively.
Insider threats can additionally undermine the fitting intentions of oversight personnel. Malicious actors with insider access can manipulate the organizational environment to align human oversight personnel’s behavior with their goals. One method is positive punishment, such as establishing a “culture of fear” that discourages human oversight personnel in engaging in behavior that is aligned with their oversight duties. Another method is positive reinforcement, where incentive structures reward behaviors deemed advantageous by the malicious actor.
4 Hardening Human Oversight
Table 2: Hardening strategies are designed to mitigate the risks of (cyber)attacks on human oversight by malicious actors. Each strategy is mapped to the specific attack vectors it protects against.
Hardening strategy Targeted attack vector
Network management (D)DoS attacks
Intrusion Detection System (D)DoS attacks, MitM attacks, malware, insider threats
Encryption MitM attacks
Transparency All attack vectors
Governance principles and audit frameworks Vulnerability Exploitation, insider threats
Red Teaming Poisoning, adversarial, and explainability attacks, vulnerability exploitation
Training of human oversight personnel Social engineering, coercion, bribery
There are various hardening strategies to minimize the risk of malicious actors exploiting attack vectors. The following list of hardening strategies draws on cybersecurity research and practice. Similar to the list of attack vectors, this list is not exhaustive but is intended to inspire approaches for designing secure human oversight. Table 2 provides an overview of mitigation strategies and indicates which attack vectors they can address in the context of human oversight.
4.1 Network Management
As discussed in section 3.4, (D)DoS attacks overwhelm a network by generating excessive traffic. Defending against these attacks requires a multi-step approach [87, 88]. The first step is to deploy monitoring systems that analyze network traffic and trigger alerts if a potential (D)DoS attack is detected. The second step involves methods to distinguish between benevolent traffic (typically generated by humans) and malicious traffic (often produced by bots). Finally, filtering techniques can mitigate (D)DoS attacks by restricting or limiting what users – humans or bots – can do within the network.
Hardening strategy: Establish robust network management procedures within the human oversight architecture.
4.2 Intrusion Detection System
Intrusion detection involves the active monitoring of computers and networks to identify attacks and misuse. An intrusion detection system (IDS) typically combines multiple components and methods to perform this monitoring. The core technical components of an IDS include network or host sensors (depending on the IDS type), a database, a management module, and an analysis module. Host and network sensors monitor the system and collect the necessary information needed for the detection of attacks. The database needs to ensure that the data collected by sensors is stored with a sufficient degree of detail and for a sufficient amount of time. The management module provides all functionalities needed for configuring and calibrating the IDS. The analysis module provides all functionalities necessary for analyzing stored data and performing incident detection. A well-designed and maintained IDS can significantly enhance the cybersecurity of IT systems [89, 59]. Intrusion detection systems may counter multiple attack vectors, including (D)DoS attacks, MitM attacks, malware, and insider threats.
Hardening strategy: Organizations should implement intrusion detection systems to protect their networks broadly and specifically to safeguard human oversight operations and the tools used by oversight personnel from malicious actors.
4.3 Encryption
Encrypted communication is the standard for data traffic on the internet. However, non-encrypted or weakly encrypted data traffic is still commonly used, particularly for communications that do not involve direct interaction between a user and a website. Such vulnerable traffic enables malicious actors to carry out Man-in-the-Middle (MitM) attacks relatively easily, often without requiring technical expertise or expensive software and hardware. To ensure adequate security, modern cryptographic methods with strong security guarantees must be employed to encrypt data traffic [90, 91].
Hardening strategy: All communication between human oversight personnel and the AI system must be encrypted. When personnel and the AI system are not co-located, using a virtual private network to protect data traffic is recommended.
4.4 Transparency
One major challenge in securing AI systems is their complexity. Enhancing transparency of both the system and its communication with human oversight personnel can improve cybersecurity by facilitating the detection of attacks. While transparency itself is not typically considered a cybersecurity measure, it provides access to key security-relevant information, such as implemented components, training data, system functionality, development processes, and dependencies. When combined with other cybersecurity tools, such as vulnerability management, the information made available through transparency increases the likelihood of detecting attacks. Additionally, transparency supports auditing and compliance attestations by making security-relevant information easily accessible and enables model tracking. Overall, ensuring transparency can also help human oversight personnel’s response time, as decision-relevant information becomes more readily available. An appropriate level of transparency can help to counter all of the attack vectors mentioned in this work.
Hardening strategy: Promote transparency to enhance attack detection by making security-relevant information accessible and making potential manipulations of the system easier to identify.
4.5 Governance Principles and Audit Frameworks
Governance principles and audit frameworks are essential for effectively managing security risks in highly automated systems. Governance principles provide the fundamental guidelines, rules, and processes that define how an organization makes decisions and ensures accountability. They establish the framework for achieving strategic goals while minimizing ethical, legal, and operational risks. In the context of AI, these principles specify who is responsible for risk assessment, how transparency is ensured, and which standards apply to data security.
A prominent example of implementing these principles is the AAA Framework, which stands for Prospective Assessments, Audit Trail, and Enforceable Compliance. These three pillars offer a comprehensive approach to strengthening system integrity and human oversight. Prospective Assessments, conducted before system implementation, identify and mitigate potential risks early, incorporating the perspectives of affected individuals and groups [66]. A robust Audit Trail, supported by high-quality data and system transparency, is essential for analyzing errors and assigning responsibility. Enforceable Compliance ensures that legal requirements are actionable, with independent audits verifying compliance and enhancing confidence in system integrity. Another well-established framework is the information security management system (ISMS) specified by ISO/IEC 27001 [92]. An ISMS provides the technical and organizational foundation to ensure integrity and reliability of information such as data, commands, and system statuses. Governance principles and audit frameworks are a countermeasure against vulnerability exploitations and insider threats.
Hardening strategy: Organizations need to establish governance principles and audit frameworks after established standards like the AAA framework and ISO/IEC 27001. These principles and frameworks have to include all components of human oversight that are implemented within an organization.
4.6 Red Teaming
The term red team refers to an independent group that investigates an organization’s IT systems to identify vulnerabilities and security flaws [93]. The red team simulates the behavior of malicious actors, employing all methods that real attackers might use. Its goal is to uncover as many exploitable security issues within the tested system as possible. Red teaming can help counter several different attack vectors, including poisoning attacks, adversarial attacks, explainability attacks, and vulnerability exploitation.
Hardening strategy: Organizations should conduct regular red teaming exercises to detect and address security flaws within their human oversight processes.
4.7 Training of Human Oversight Personnel
Training aims to strengthen the resilience of human oversight personnel against attacks [94]. Training initiatives can help personnel to recognize, resist, and properly respond to manipulation attempts. Training may include social engineering awareness, teaching personnel to identify common manipulation tactics such as phishing or impersonation of authority figures. Practical exercises, such as simulated phishing campaigns or role-play scenarios, reinforce recognition and proper reporting behaviors.
Training can also address resilience against coercion, preparing personnel to respond to threats or undue pressure. This includes protocols for escalating coercion attempts, access to secure whistleblowing channels, and preparation to maintain effective human oversight behavior under stress. Anti-bribery training can teach personnel to recognize undue influence. Clear organizational policies defining conflicts of interest and mandatory disclosure mechanisms further reduce the risk of misaligned intentions. Training of human oversight personal counters social engineering, coercion, and bribery.
Hardening strategy: Organizations should implement recurring, context-specific training programs for human oversight personnel. By institutionalizing such training and embedding it in the organization’s security culture, human oversight personnel can develop the awareness, self-efficacy, and procedural support needed to counteract manipulation attempts, thereby contributing to secure human oversight.
5 Discussion
Research on human oversight of AI has primarily focused on increasing its effectiveness [1, 2, 6, 4]. In this paper, we highlight the importance of complementing this focus with a security-related perspective on human oversight. This perspective is critical because various attack vectors can undermine human oversight effectiveness, a risk that will grow as human oversight becomes more widely implemented as part of the safety, security, and accountability architecture of AI operations. Drawing on cybersecurity research [33, 44], we have introduced various attack vectors targeting human oversight, along with corresponding hardening strategies, to support the development of secure human oversight of AI.
Understanding and designing the technical factors, selecting and training human oversight personnel, and shaping working conditions to foster effective human oversight remains an ongoing challenge for research, practice, and AI governance [1, 2, 95]. Our analyses indicate that this challenge will intensify as human oversight becomes standard practice in high-risk AI use cases. Malicious actors targeting the attack surface of human oversight can compromise the AI system, the communication between humans and AI, or the oversight personnel themselves, aiming to undermine epistemic access, causal power, self-control, and fitting intentions as the core requirements for effective oversight [1]. Consequently, such attacks may increase risks in AI operations by reducing the effectiveness of oversight, disabling oversight entirely, or manipulating oversight personnel’s intentions to exacerbate risks. The attack vector list we present is not exhaustive; future research should explore novel attack vectors specific to human oversight. A potential future scenario involves the deployment of malicious AI agents designed to deceive oversight [70, 69]. Once an AI system models the oversight mechanism, it may attempt to circumvent the restrictions imposed by human oversight, further expanding the attack surface and associated risks.
Given these risks, it is essential to consider how to enhance the security of human oversight. To this end, we presented various hardening strategies drawn from the broader cybersecurity literature [52, 53, 54]. These strategies must be actively applied and adapted to the context of human oversight of AI. Our paper contextualizes these strategies for human oversight, representing a crucial first step toward securing oversight architectures against malicious attacks. Future research needs to evaluate the effectiveness of these hardening strategies and could develop additional strategies tailored specifically to human oversight. Beyond technical hardening, future work could also explore the application of amplified oversight. The core principle of amplified oversight is to employ AI itself to systematically reduce complexity to a level that enables human oversight personnel to fulfill their duties effectively [96]. In future, such approaches to support human oversight may even exceed human capabilities and may become a further viable target for malicious actors to attack.
While this paper focuses on malicious actors targeting human oversight, not all risks emerge from malicious actors. Additional risks can for example arise from counterproductive behavior within organizations [97]. Counterproductive behavior refers to intentional or unintentional actions by employees that go against organizational goal, such as neglecting oversight duties, resisting established procedures, or even deliberately undermining AI operations, which can erode the effectiveness of human oversight. Such behavior may stem from factors like dissatisfaction with working conditions, misaligned incentives, organizational conflict, or lack of recognition [97]. Such counterproductive behavior inside organizations may need closer examination in the future to ensure secure human oversight. Such behavior can undermine the core requirements of effective oversight and may promote active damage to the organization or passive under-performance, thereby introducing risks to human oversight of AI. Although these behaviors are not attacks by malicious actors, they may make attacks by external actors more effective, for example, if an attacker targets a dissatisfied employee. Future works could shed light on those risk factors and suitable hardening strategies such as organizational structures, incentive systems, and thoughtful work design [29]. Well-designed security concepts, policies and work conditions can all together mitigate security risks and prevent counterproductive behavior by ensuring that oversight personnel’s goals remain aligned with organizational objectives.
6 Conclusion
Research on human oversight has primarily focused on its effectiveness – and rightfully so, given the complexity of designing oversight processes that genuinely mitigate risks in AI operations [1, 3, 2]. Adding to this complexity, it will become increasingly important to design for secure human oversight as oversight becomes an integral component of the safety, security, and accountability infrastructure of AI operations. We hope this paper serves as a starting point for research that analyzes the attack surface of human oversight and develops effective hardening strategies to prevent malicious actors from exploiting oversight as a gateway to compromise AI safety.
Acknowledgments
The work of Markus Langer is partially funded by the the Daimler and Benz Foundation as part of the project TITAN – Technologische Intelligenz zur Transformation, Automatisierung und Nutzerorientierung des Justizsystems (grant no. 45-06/24). The work of Kevin Baum is partially funded by the German Federal Ministry of Education and Research (BMBF) as part of the project MAC-MERLin (Grant Agreement No. 01IW24007), and the European Regional Development Fund (ERDF) as well as the German Federal State of Saarland within the scope of the Project ToCERTAIN. In addition, the work of Markus Langer and Kevin Baum is partially funded by the DFG grant 389792660 as part of TRR 248 CPEC — Center for Perspicuous Computing. This work has benefitted from the sixth and last authors’ participation in Dagstuhl Seminar Dagstuhl Seminar 25272 "Challenges of Human Oversight: Achieving Human Control of AI-Based Systems".

### v2

@@
-Secure Human Oversight of AI: Exploring the Attack Surface of Human Oversight
+Secure human oversight of AI: Threat modeling in a socio-technical context
 Abstract
@@
 We argue that human oversight creates a new attack surface within the safety, security, and accountability architecture of AI operations.
-Drawing on cybersecurity perspectives, we analyze attack vectors that threaten the requirements of effective human oversight, thereby undermining the safety of AI operations.
-Such attacks may target the AI system, its communication with oversight personnel, or the personnel themselves.
-We then outline hardening strategies to mitigate these risks.
-Our contributions are: (1) introducing a security perspective on human oversight, and (2) providing an overview of attack vectors and hardening strategies to enable secure human oversight of AI.
+Drawing on cybersecurity perspectives, we model human oversight as an IT application for the purpose of systematic threat modeling of the human oversight process.
+Threat modeling allows us to identify security risks within human oversight and points towards possible mitigation strategies.
+Our contributions are: (1) introducing a security perspective on human oversight, (2) offering researchers and practitioners guidance on how to approach their human oversight applications from a security point of view, and (3) providing a systematic overview of attack vectors and hardening strategies to enable secure human oversight of AI.
 Keywords Human oversight $\cdot$ cybersecurity $\cdot$ attack vectors $\cdot$ hardening strategies $\cdot$ AI governance
 1 Introduction
-Figure 1: The human oversight attack surface encompasses the AI system’s technical infrastructure, the human oversight personnel, and the communication channels between the two sides.
-Attacks on human oversight aim to undermine oversight personnel’s epistemic access, causal power, self-control, and fitting intentions, the four requirements of effective human oversight [1].
-Malicious actors can utilize different attack vectors to undermine one or several of these requirements (see Table 1).
-In order to protect human oversight from malicious actors, various hardening strategies need to be considered and deployed to counteract the threats posed by the attack vectors (see Table 2).
-A well-designed combination of hardening strategies can shield against different attack vectors and, thus assure secure human oversight.
 Human oversight of artificial intelligence (AI) is often considered a strategy to address risks associated with the use of AI in high-risk contexts [25, 45, 67].
@@
 Human oversight is also a key concept in AI governance and ethics [45, 25] and a central requirement for the implementation of high-risk AI systems in emerging AI regulation, such as Article 14 of the European AI Act [18, 45].
-However, the idea of relying on humans to oversee AI in order to mitigate operational risks also faces criticism and is expected to encounter substantial challenges in practice [7, 2, 1].
+However, relying on humans to oversee AI also faces criticism and substantial challenges in practice [34, 25, 72].
 The most prominent criticisms and challenges include doubts about whether humans can effectively oversee AI systems [25], the difficulty of defining what constitutes effective human oversight [72], clarifying who should provide oversight and when it should occur [18], specifying which risks and unwanted system behaviors oversight personnel should detect [41], and determining how to design and evaluate effective human oversight [72].
@@
 In this sense, the human oversight architecture creates a new attack surface for targeting AI operations in high-risk contexts such as critical infrastructure, public administration, or medicine.
-A security-focused perspective on human oversight is crucial, given the central role oversight already plays, the expectations placed on it in high-risk contexts, and its growing importance as a regulatory requirement in AI governance, such as under the EU AI Act.
-If we fail to assess and address the security challenges of human oversight, we risk overlooking potential attack vectors in high-risk contexts; and without awareness of these vectors, we will not be able to design effective hardening strategies.
+A security-focused perspective on human oversight is crucial, given the central role human oversight already plays, the expectations placed on it in high-risk contexts, and its growing importance as a regulatory requirement in AI governance, such as under the EU AI Act.
+If we fail to assess and address the security challenges of human oversight, we risk overlooking potential attack vectors in high-risk contexts; and without awareness of these vectors, we will not be able to design effective countermeasures and hardening strategies.
+Ultimately, neglecting security aspects of human oversight can undermine its main objective—risk mitigation—or even increase risks by introducing additional attack vectors.
 In this paper, we introduce a security-related perspective on human oversight of AI.
-Specifically, drawing on cybersecurity research, we outline possible attack vectors that threaten the key requirements of effective human oversight (i.e., epistemic access, self-control, causal power, and fitting intentions; [1]).
-These attacks may target the AI system itself, the communication between the AI system and oversight personnel, or the personnel directly (see Figure 1).
-We also draw on the cybersecurity literature to discuss hardening strategies that can mitigate the risks posed by such attacks.
-This paper makes two key contributions.
+Human oversight is best understood as a socio-technical system that can be instantiated in many different ways.
+For the purposes of threat modeling, we model it as an IT application and draw on established cybersecurity methods to analyze attack surfaces and vulnerabilities.
+This enables us to identify security threats to the key requirements of effective human oversight (i.e., epistemic access, self-control, causal power, and fitting intentions; [72]).
+Threat modeling further provides insights into applicable countermeasures and hardening strategies to mitigate the identified cybersecurity risks.
+This paper makes three key contributions.
 First, we introduce a security-related perspective on human oversight of AI.
-This perspective is critically missing from current research, even though human oversight is becoming an integral component of the safety, security, and accountability architecture of AI operations, and thus an increasingly viable target for malicious actors.
-Second, we provide an overview of possible attack vectors and hardening strategies contextualized for human oversight of AI.
-We hope this contribution supports AI governance and system design in ensuring not only effective but also secure human oversight of AI.
+This perspective is critically missing from research, even though human oversight is becoming an integral component of the safety, security, and accountability architecture of AI operations, and thus an increasingly viable target for malicious actors.
+Second, we provide guidance on how to perform threat modeling in socio-technical contexts and identify cybersecurity threats of human oversight with possible countermeasures.
+Thereby, this work provides a starting point for future threat modeling of implementations of human oversight.
+Third, we provide a list of security vulnerabilities and hardening strategies that are applicable to human oversight applications in general and should always be considered by practitioners.
+We hope these contributions support AI governance and system design in ensuring not only effective but also secure human oversight of AI.
 2 Related Work
@@
 Research discusses the nature of human oversight of AI, including the what, when, and who of oversight.
-Regarding the what, many forms of human-AI collaboration in principle involve some degree of human oversight.
+Regarding the what, many forms of human-AI collaboration involve human oversight.
 Across a wide range of human-AI interaction use cases, humans monitor AI system outputs, check their accuracy, request explanations, adapt outputs, intervene when systems malfunction, or take over manual control in the event of failures [40, 10, 81, 17, 47, 73, 46, 13].
-In a more targeted sense, and following how emerging research has tentatively defined the concept [1], human oversight specifically refers to humans mitigating risks in AI operations [16, 6].
-More concretely, oversight personnel are tasked with monitoring AI systems, detecting failures and malfunctions, identifying unsafe states, flagging inaccurate outputs, and recognizing outputs that may pose risks to fundamental human rights, such as discriminatory outputs [1, 5, 6].
-Once such issues are detected, oversight personnel are expected to act in ways that effectively mitigate emerging risks [5, 1].
+In a more targeted sense, and following how emerging research has defined the concept [72], human oversight specifically refers to humans mitigating risks in AI operations [50, 18].
+More concretely, oversight personnel should monitor AI systems, detect failures and malfunctions, identify unsafe states, flag inaccurate outputs, and recognize outputs that may pose risks to fundamental human rights, such as discriminatory outputs [72, 41, 18].
+Once such issues are detected, oversight personnel are expected to act in ways that effectively mitigate risks [41, 72].
 This can involve taking over manual control, overwriting inaccurate or inadequate outputs [14], or stopping AI operations and triggering intervention by other human stakeholders, such as developers who can adapt the system.
@@
 For example, a patient using an AI chatbot for mental health support [76] would be the user, while oversight personnel monitor the patient–AI interaction to detect risky situations (e.g., if the patient discusses suicidal behavior with the chatbot).
-Regarding the when of human oversight, human involvement is important at every stage of AI development: during system design, at run time, and when inspecting errors and failures in hindsight [4, 19, 20].
+Regarding the when, human involvement is important at every stage of AI development: during system design, at run time, and when inspecting errors and failures in hindsight [67, 68, 24].
 In a more targeted sense, human oversight usually refers to personnel overseeing AI systems during run time [72, 45].
-This can involve dynamic and time-critical situations, such as humans overseeing autonomous vehicles [17, 21], but also less dynamic contexts, such as AI in judicial decision-making or hiring, where there is more time to deliberate about the accuracy or adequacy of AI outputs [5, 22].
+This ranges from dynamic, time-critical situations like autonomous vehicle oversight [14, 39] to less time-pressured contexts like judicial AI [41, 33].
 Regarding the who of human oversight, any human with the skills necessary to effectively oversee an AI system could serve as oversight personnel [18, 78].
@@
 [72] identify four key aspects: epistemic access, causal power, self-control, and fitting intentions.
-As we discuss possible attack vectors, we consider how successful attacks might undermine these requirements.
 Epistemic access means that oversight personnel possess sufficient knowledge about the AI system and understand the consequences of their interventions.
 Without it, they may fail to detect system failures or inaccurate outputs.
-Causal power refers to the ability to intervene in ways that actually influence AI processes and outputs.
-Without causal power, oversight personnel may be unable to overwrite or correct system outputs.
+Causal power refers to the ability to intervene in ways that influence AI processes and outputs.
+Without it, oversight personnel may be unable to overwrite or correct system outputs.
 Self-control requires oversight personnel to maintain agency over their own actions.
@@
 2.2 Attack Vectors and Hardening Strategies in Cybersecurity
-With the implementation of AI systems across diverse fields of application, ensuring the security of deployed solutions has become increasingly urgent.
-Achieving this requires an in-depth understanding of specific attack vectors and the attack surface within each application domain [30].
+With the implementation of AI systems across diverse context, ensuring the security of deployment has become increasingly important.
+Enabling systematic threat modeling for AI-based applications requires an in-depth understanding of specific attack vectors and the attack surface within an application domain [74].
 Traditionally, an attack vector is defined as a method or pathway that a malicious actor uses to gain access to a network or system.
@@
 We use the term attack surface to describe the total set of points where a malicious actor could attempt to enter, exploit, or otherwise attack digital and AI infrastructures [74].
-For example, research has systematically analyzed attack vectors in autonomous driving systems and defined the corresponding attack surface [31].
-Other work has examined the attack surface in federated learning workflows [32].
-For human oversight, however, such systematic investigations of the attack surface are still missing.
-Relevant related work can be found in the field of AI cybersecurity, which examines attack vectors that explicitly target AI systems.
+While there is research that systematically analyses AI-based systems from a cybersecurity point of view [22, 60], such systematic investigations are still missing for human oversight.
+Related work can be found in the field of AI cybersecurity, which examines attack vectors that explicitly target AI systems.
 For instance, poisoning attacks manipulate training data to introduce backdoors into AI models that can later be exploited [65, 26, 32].
@@
 Such attacks manipulate explanations to obscure the real reasons behind a model’s decision [70, 1, 69].
-Because human oversight extends beyond an isolated AI system, additional relevant research comes from traditional cybersecurity.
-Distributed denial-of-service (DDoS) attacks flood networks with traffic, making them unresponsive or unavailable [42, 43].
-Man-in-the-middle attacks intercept and manipulate data traffic between two endpoints [44].
-Another common vector is malware, which – due to its diversity – can be deployed against nearly any digital system [45, 46].
-Particularly relevant for defining the attack surface are attacks on the human factor in cybersecurity.
-Social engineering exploits human vulnerabilities to extract sensitive information or gain access privileges [47, 48].
-Two additional but in cybersecurity research less discussed vectors are coercion and bribery, which can likewise be used to gain sensitive information or access privileges [49, 50].
-Finally, insider threats exploit individuals inside an organization (e.g., disgruntled employees) to carry out attacks [51].
-With the growing threat of cyberattacks, every system that incorporates digital components must be analyzed for potential attack vectors.
-Our work contributes to this literature by contextualizing attack vectors for the specifics of human oversight of AI.
-In response to the many ways digital systems can be attacked, cybersecurity research has developed hardening strategies as methods designed to secure systems against attacks [52, 53, 54].
+However, AI system are still IT systems at the core.
+This means that traditional attack vectors, like distributed denial-of-services (DDOS) [43, 58], man-in-the-middle (MITM) [15], or social engineering attacks [5, 16], have to be considered for AI systems as well.
+Furthermore, human-centered attack vectors that are usually underexplored in cybersecurity, e.g., coercion, bribery, and insider threats like disgruntled employees [28, 75, 48], are of special importance for human oversight.
+In response to the many ways digital systems can be attacked, cybersecurity research has developed hardening strategies as methods designed to secure AI systems against attacks [62, 61, 9].
 One line of work establishes methods for securing data traffic in computer networks.
@@
 Research has explored how to improve system security by training users and raising awareness of risks [23, 55].
-In this paper, we extend the literature by contextualizing hardening strategies specifically for securing human oversight of AI against malicious attacks.
-3 The Human Oversight Attack Surface
+In this paper, we extend the literature by contextualizing attack vectors and hardening strategies specifically for securing human oversight of AI and show how to use this knowledge for systematic threat modeling of human oversight.
+3 Threat modeling of human oversight
 Human oversight can serve as part of the safety, security, and accountability architecture of AI systems, as oversight personnel are expected to detect and mitigate risks.
 In addition, oversight personnel may be required to actively confirm that an AI system is in a safe state.
-Human oversight may become even more important in AI operations if it evolves into a central requirement of AI governance and regulation in high-risk contexts [2, 6, 3, 1].
-With a more wide-spread implementation, human oversight will also become an increasingly attractive target for malicious actors seeking to corrupt or disrupt AI-based operations.
+Human oversight may become even more important in AI operations if it evolves into a requirement of AI governance and regulation in high-risk contexts [25, 18, 45, 72].
+With a more wide-spread implementation, human oversight will also become an increasingly viable target for malicious actors seeking to corrupt or disrupt AI-based operations.
 Such actors could be cybercriminals targeting high-risk AI systems for ransom or foreign governments attacking a country’s critical infrastructure in hybrid or open conflict.
 Furthermore, sufficiently powerful AI agents that are successfully attacked could themselves become malicious actors by attempting to evade oversight [7, 51].
-Considering these threats, designing human oversight requires not only ensuring effective oversight [1] but also ensuring secure oversight.
-Designing for secure human oversight demands an understanding of its attack surface.
-The following list of attack vectors inspired by the broader cybersecurity literature is not exhaustive but illustrates the breadth of the human oversight attack surface.
-It outlines attacks that target the four core requirements of effective human oversight: epistemic access, self-control, causal power, and fitting intentions.
-Table 1 summarizes these attack vectors and their relation to the requirements for effective human oversight.
-Table 1: Attack vectors that target the effectiveness of human oversight. For each attack vector, we indicate the specific requirement for effective human oversight that the attack seeks to undermine.
-Attack vectors Affected requirements of effective human oversight
-Poisoning attacks Epistemic access, self-control
-Adversarial attacks Epistemic access
-Explainability attacks Epistemic access
-(Distributed) denial of service attacks Causal power
-Man-in-the-middle attacks Epistemic access, causal power
-Vulnerability exploitation Epistemic access, causal power
-Malware Epistemic access, causal power
-Social engineering Epistemic access, self-control
-Coercion Self-control
-Bribery Fitting intentions
-Insider threats Epistemic access, causal power, fitting intentions
-3.1 Poisoning Attacks
-Poisoning attacks target the core functionality of AI models, namely the association of patterns in data with outcomes (e.g., real-valued numbers in regression, class probabilities in classification, or new text/image data in generative AI) [71].
-These attacks exploit the fact that modern AI models are often trained on large volumes of (semi-)automatically collected data, for example through web crawlers.
-Malicious actors can insert triggers, or predefined patterns, into such datasets.
-When models are trained on compromised data, they may consistently exhibit malicious behaviors whenever a trigger is presented [72].
-Poisoning attacks can compromise both the epistemic access and the self-control of oversight personnel.
-They undermine epistemic access because poisoned samples containing hidden triggers cause AI responses that deviate from normal outputs.
-A carefully crafted trigger, such as an imperceptible pixel pattern in images [73], can prevent oversight personnel from recognizing the behavioral shift.
-Even when oversight personnel detect that something is amiss, hidden triggers may still make it difficult to understand the malicious behavior.
-Poisoning attacks can also erode self-control.
-If triggered outputs cause the oversight interface to generate excessive micro-notifications or confirmation requests, malicious actors can exploit this to induce fatigue by overwhelming oversight personnel with trivial but time- and energy-consuming tasks.
-This not only reduces oversight efficiency by consuming resources but also increases the likelihood of errors once fatigue reaches a critical level, thereby undermining the security of human–AI interaction.
-3.2 Adversarial Attacks
-Adversarial attacks aim to induce misclassifications in the outputs of an AI model [71].
-By feeding the system with adversarial examples, attackers seek either to fool the model or to reduce its accuracy and performance.
-These adversarial examples typically resemble normal inputs but include subtle perturbations that may not be perceptible to oversight personnel, leading to misclassifications aligned with the attacker’s goals [72].
-Adversarial attacks primarily compromise the epistemic access of oversight personnel.
-A crucial part of understanding AI behavior is knowing how input patterns are mapped to outputs (e.g., which structures or features in an input image correspond to specific output classes).
-Without this knowledge, reliably determining whether the system behaves acceptably becomes exceedingly difficult.
-Adversarial perturbations obfuscate the relationship between inputs and outputs, reducing epistemic access to the model’s true behavior.
-As a result, oversight personnel may fail to detect unwanted model behavior or even system failures.
-3.3 Explainability Attacks
-Attacks on explainability specifically target explainability and interpretability approaches and tools.
-These attacks manipulate or deceive the evidence for a model’s reasoning.
-A key characteristic of such attacks is their ability to alter the explanation while keeping the model’s output unchanged.
-This means that the AI system may still produce the same output, but the explanation for that output is inaccurate or misleading [74].
-Explainability attacks primarily undermine the epistemic access of oversight personnel.
-Misleading or inaccurate explanations prevent personnel from understanding the AI system’s true behavior.
-Attackers can exploit these attacks to fabricate behavior that serves their goals.
-Even with extensive querying, oversight personnel may be unable to detect such manipulations [75].
-Moreover, a mismatch between outputs and explanations adversely affects personnel who rely on explanations to identify problematic AI behavior [25, 76].
-These attacks can also induce a state of deceptive epistemic access, where the system appears transparent but provides inaccurate explanations.
-This scenario can be more dangerous than a “black box” system because it fosters unjustified trust and makes detecting incorrect outputs more difficult.
-Consequently, a key focus of explainability for human oversight must be ensuring the fidelity and robustness of explanations [77, 78].
-3.4 (Distributed) Denial of Service Attacks
-Denial of service (DoS) attacks refer to methods that render a system inaccessible [79].
-The core principle of DoS attacks is to overwhelm a target system with more requests than it can handle, causing the system to become nonfunctional.
-Various techniques exist to perform a DoS attack, such as exploiting the TCP/IP protocol (SYN flooding), using the ping functionality (ping flooding), or targeting mail servers (mail flooding).
-A variation of DoS attacks is the distributed denial of service (DDoS) attack, which distributes requests across multiple systems rather than originating from a single source [80].
-(Distributed) DoS attacks can undermine the causal power of oversight personnel.
-Modern computing systems are increasingly implemented as network-based, distributed systems—a standard likely to apply in human oversight settings, where the AI system runs on a cloud-based or on-premise server while oversight personnel interact with it via network protocols.
-(D)DoS attacks can make either the AI system or the human oversight interface inaccessible.
-This threat is especially severe for cloud-based AI systems: because the AI system and oversight personnel are physically separated, malicious actors can effectively disable oversight through (D)DoS attacks while the AI system continues to operate in an unsafe state or produces inaccurate outputs.
-3.5 Man-in-the-Middle Attacks
-An attack vector that can be used in computer networks is the Man-in-the-Middle (MitM) attack [44].
-When successful, this strategy allows malicious actors to intercept and control data traffic between two or more systems, granting them access to the exchanged information and enabling them to manipulate it.
-Additionally, attackers can conceal their presence, tricking the involved systems into believing they are communicating securely with each other [44].
-Various methods exist to carry out a MitM attack: ARP spoofing can be used when the attacker and target are on the same local area network (LAN), snarfing can be applied in WLAN-based attacks, and domain name system (DNS) cache poisoning is another possible approach.
-MitM attacks can compromise the epistemic access and causal power of oversight personnel, depending on which direction the manipulated information travels.
-If the attacker intercepts and alters information from the AI system to the oversight personnel, epistemic access is undermined: the true behavior, functionality, or outputs of the AI system may remain hidden.
-If the manipulated information flows from the human to the AI system, causal power is compromised.
-For example, if an attacker modifies commands issued by oversight personnel, they lose the ability to interact effectively with the AI system and cannot intervene in AI processes.
-This situation is particularly dangerous because the attacker can send manipulated responses back to the personnel, creating an illusion of control and misleading them into believing they are still effectively supervising the AI system.
-3.6 Vulnerability Exploitation
-Vulnerability exploitation refers to utilizing existing flaws in hardware and software for attacks, such as remote code execution or exfiltration of sensitive data.
-Since modern digital systems are complex combinations of hardware and software components and their respective supply chains, flaws can accumulate quickly and pose a significant threat.
-One of the most dangerous vulnerability exploitation is the zero-day exploit [81].
-This term refers to the exploitation of a specific vulnerability within a hardware or software component that is unknown even to the manufacturer of the component.
-This gives the attacker the time advantage of day zero, before the manufacturer can detect, identify and fix the vulnerability.
-Attacks based on vulnerability exploitation can target epistemic access or causal power.
-Attackers can use the vulnerabilities to get unauthorized access to the system and in this way manipulate the system output which the human oversight person uses to detect deviations in system behavior.
-If the attacker can obfuscate their actions from the human oversight personnel, the personnel are prevented from understanding the real functioning of the AI system.
-Similar effects can be achieved by manipulating other information that human oversight personnel uses to assess the AI system.
-Causal power is affected by vulnerability exploitation if attackers prevent human oversight personnel from appropriately communicating with the AI system.
-This can be achieved by dropping commands send by the human oversight personnel or disabling crucial components of the human oversight software.
-Depending on the exploitation, attackers might be able to disguise the attack.
-For example, and similar to the example for MitM attacks, they could drop commands by human oversight personnel but fake a response by the AI system that suggests the command was successfully executed.
-3.7 Malware
-Malware is software that executes harmful functionality on an IT system, often without the user’s knowledge.
-Its potential targets and effects are highly diverse.
-Ransomware encrypts the file system and typically demands a ransom for decryption.
-Spyware extracts information from the infected system.
-These are just two of the most common types, but malware encompasses a wide range of threats designed for various purposes [82, 45].
-Due to the diversity of malware functionality, it can compromise epistemic access and causal power of oversight personnel.
-Regarding epistemic access, malware can alter the core functionality of an AI system.
-For example, it could insert system prompts into deployed large language models (LLMs), significantly changing their behavior [83].
-Malware can also intercept or manipulate data traffic from the AI system to oversight personnel, or modify, encrypt, or obfuscate system logs.
-These actions prevent personnel from understanding the system’s true functionality and therefore undermine epistemic access.
-Causal power can be compromised if malware intercepts or manipulates data traveling from oversight personnel to the AI system.
-For instance, if malware prevents personnel from interacting with the AI system or locks them out of controlling its core functions, they lose the ability to intervene in system processes or outputs.
-Similarly, if malware secretly alters commands issued by oversight personnel, it further undermines their causal power.
-3.8 Social Engineering
-Social engineering refers to attacks that exploit human characteristics [84].
-These attacks leverage traits such as fear, respect for authority, or the willingness to help, tricking victims into revealing personal information, disabling security measures, or installing malware.
-The widespread use of digital communication channels has provided malicious actors with effective and scalable tools for conducting social engineering attacks [85].
-Social engineering can compromise epistemic access and self-control of oversight personnel.
-If an attacker impersonates individuals with technical expertise or authority (e.g., trusted colleagues, management, external experts) they can distort personnel’s understanding of relevant information needed to oversee AI systems.
-For example, an attacker might pretend to be a technician from the AI provider, convincing human oversight personnel that certain graphical user interface notifications indicate a bug.
-This could lead personnel to ignore important alerts or install malware disguised as a hotfix.
-Similarly, attackers impersonating management may convince oversight personnel that specific system notifications are non-critical, framing them as part of corporate processes.
-In both cases, social engineering feeds false information about the AI system’s functionality to oversight personnel, thereby undermining epistemic access.
-Social engineering can also erode self-control.
-By impersonating senior colleagues or management, attackers can exploit perceived authority to pressure oversight personnel into taking specific actions.
-If the attack is credible, personnel might follow instructions against their better judgment, believing the directives come from a legitimate authority.
-3.9 Coercion
-Coercion is a broad concept that extends beyond IT security and involves using force, often through threats or violations, to influence a party.
-In the context of human oversight, coercion refers to pressure or force that alters the intended actions of oversight personnel.
-Coercion undermines the self-control of oversight personnel.
-By influencing oversight personnel to act according to the instructions of a malicious actor rather than their own judgment, coercion renders human oversight ineffective.
-Even when oversight personnel detect a need for intervention, they may be unable to act according to their assessment.
-In extreme cases, a malicious actor could threaten the oversight personnel to gain control over their actions, effectively nullifying the intended benefits of human oversight.
-3.10 Bribery
-Bribery refers to the act of persuading a person, organization, or authority to act in accordance with the goals of a malicious actor, typically by offering monetary incentives, status, or similar benefits.
-Bribery undermines the fitting intentions of human oversight personnel.
-Usually, the intentions of human oversight personnel should align with fulfilling their oversight duties, such as detecting inaccurate outputs or system failures [1].
-A successful bribe can misalign these intentions, causing personnel to act in accordance with the malicious actor’s goals instead.
-Such goals might include damaging critical infrastructure or compromising the functioning of AI in high-risk applications, such as in hospitals.
-3.11 Insider Threats
-Insider threats encompass attack scenarios in which malicious actors infiltrate the organization they intend to compromise [86].
-Such attacks share the common characteristic that the attacker either has direct access themselves or exploits an employee’s access rights, whether voluntarily or accidentally.
-Two pathways to insider threats are discussed under the sections on coercion (see 3.9) and bribery (see 3.10).
-Another common pathway is the exploitation of disgruntled employees.
-Insider threats can undermine epistemic access, causal power, and fitting intentions of oversight personnel.
-They can erode epistemic access by altering or falsifying monitoring data from the AI system.
-Since human oversight personnel rely on this data to accurately assess AI behavior and intervene when necessary, this prevents them from understanding the system.
-Insider threats can also obfuscate AI system documentation by deleting or modifying key information, further impeding epistemic access.
-Insider threats can also compromise the causal power of oversight personnel.
-For example, possible sequences of actions that can be taken as a response to detecting problematic system behavior can be ambiguous and the effectiveness of different sequences of actions can differ vastly.
-Attackers may be able to change organizational oversight policies on how to deal with inaccurate outputs or system malfunctions.
-In the worst case, such adapted policies can prevent oversight personnel from acting appropriately after detecting problematic system behavior which negatively affects causal power.
-Another example of how insider threats may undermine causal power is by achieving and thereby exploit under-staffing.
-Redirecting resource allocation away from human oversight or preventing human oversight personnel from doing their work (e.g., by force) can reduce the workforce available to human oversight.
-While this requires substantial access and intervention capabilities, it can prevent oversight personnel from exercising causal power effectively.
-Insider threats can additionally undermine the fitting intentions of oversight personnel.
-Malicious actors with insider access can manipulate the organizational environment to align human oversight personnel’s behavior with their goals.
-One method is positive punishment, such as establishing a “culture of fear” that discourages human oversight personnel in engaging in behavior that is aligned with their oversight duties.
-Another method is positive reinforcement, where incentive structures reward behaviors deemed advantageous by the malicious actor.
-4 Hardening Human Oversight
-Table 2: Hardening strategies are designed to mitigate the risks of (cyber)attacks on human oversight by malicious actors. Each strategy is mapped to the specific attack vectors it protects against.
-Hardening strategy Targeted attack vector
-Network management (D)DoS attacks
-Intrusion Detection System (D)DoS attacks, MitM attacks, malware, insider threats
-Encryption MitM attacks
-Transparency All attack vectors
-Governance principles and audit frameworks Vulnerability Exploitation, insider threats
-Red Teaming Poisoning, adversarial, and explainability attacks, vulnerability exploitation
-Training of human oversight personnel Social engineering, coercion, bribery
-There are various hardening strategies to minimize the risk of malicious actors exploiting attack vectors.
-The following list of hardening strategies draws on cybersecurity research and practice.
-Similar to the list of attack vectors, this list is not exhaustive but is intended to inspire approaches for designing secure human oversight.
-Table 2 provides an overview of mitigation strategies and indicates which attack vectors they can address in the context of human oversight.
-4.1 Network Management
-As discussed in section 3.4, (D)DoS attacks overwhelm a network by generating excessive traffic.
-Defending against these attacks requires a multi-step approach [87, 88].
-The first step is to deploy monitoring systems that analyze network traffic and trigger alerts if a potential (D)DoS attack is detected.
-The second step involves methods to distinguish between benevolent traffic (typically generated by humans) and malicious traffic (often produced by bots).
-Finally, filtering techniques can mitigate (D)DoS attacks by restricting or limiting what users – humans or bots – can do within the network.
-Hardening strategy: Establish robust network management procedures within the human oversight architecture.
-4.2 Intrusion Detection System
-Intrusion detection involves the active monitoring of computers and networks to identify attacks and misuse.
-An intrusion detection system (IDS) typically combines multiple components and methods to perform this monitoring.
+Considering these threats, designing human oversight requires not only ensuring effective oversight [72] but also ensuring secure human oversight.
+In cybersecurity, the standard approach to identify, quantify and address potential flaws and vulnerabilities within an IT system is to perform threat modeling.
+Organizations and IT-professionals can use threat modeling to examine the IT system at hand from the viewpoint of malicious actors using a structured process.
+Threat modeling consists of four steps.
+First, the scope of the tested application has to be determined, usually by creating a data flow diagram (DFD).
+Furthermore, the first step identifies entry points into the application, exit points of the application, privilege or trust boundaries, and assets that need to be protected.
+The second step involves determining threats to the tested application.
+In this work, the STRIDE [37] model is used for determining threats.
+STRIDE—an acronym for Spoofing, Tampering, Repudiation, Information disclosure, Denial of service, Elevation of privilege—is a well-established threat model.
+During the third step of threat modeling, countermeasures and potential hardening strategies are determined for identified threats.
+The fourth step consists of assessing the results of the previous steps and ensuring that no important points have been missed.
+For this manuscript, step four will be implicitly addressed in the discussion.
+This section details the decisions we took in performing threat modeling for an IT-application representation of human oversight, and how we incorporated its socio-technical context into the analysis.
+We provide a guideline for researchers and practitioners who want to perform threat modeling for their real-world implementations.
+3.1 Step 1 - The scope of human oversight
+Figure 1: Data flow diagram (DFD) of the abstracted human oversight application used for threat modeling.
+Arrows indicate the directed paths of data through the application.
+Each arrow is labeled with the type of data that is transported.
+Circles indicate processes, two concentric circles indicate multiple subprocesses, and rectangles indicate external entities.
+Privilege boundaries are drawn with red dashed lines.
+HO = Human Oversight.
+As mentioned above, the first step of any threat modeling is to collect information about the subject of the threat modeling, and thereby determine the scope of the work.
+A data flow diagram (DFD) shows different paths through an application and indicates the locations of privilege boundaries (also called trust boundaries).
+These boundaries indicate that movement along the path leads to a change of privileges, i.e., users and processes beyond the boundary have more rights to perform actions within the system like write data or run software.
+We extend the meaning of privilege boundaries to also include the right for intervention, meaning that entities with higher privilege can intervene on entities with lower privilege.
+The shape of nodes in a DFD are associated with specific meanings.
+A rectangle indicates an external entity, like a user that accesses the application.
+A circle indicates a single process.
+Two concentric circles indicate multiple subprocesses.
+Data flows are indicated by arrows, and privilege boundaries are indicated by dashed lines.
+Traditionally, DFDs can indicate data storage nodes with two parallel lines but we did not need this node type for the human oversight DFD.
+Furthermore, DFDs are hierarchical in nature.
+This means that several different levels of DFDs can be created for an application.
+Each DFD focuses on a different part of the application.
+This approach has the advantage that every isolated DFD remains relatively simple.
+We use this property to compensate for the fact that we are not looking at a concrete implementation but a more abstract concept.
+Designing a high-level DFD allows us to omit implementation details via the two types of process nodes.
+Real-world implementations of human oversight can than fill in the details of the processes with lower-level DFDs.
+3.1.1 The human oversight DFD
+Figure 1 presents our DFD of the human oversight application.
+We identified three external entities.
+The users node includes all entities that interact with the AI system, either humans or other IT systems.
+The superior node indicates structures within the organization that have greater authority than the human oversight system.
+That could be for example the management of the organization.
+The external advisory board node indicates an external entity that defines appropriate actions and goals of human oversight, similar to an ethical advisory boards.
+Next, there are two multiple processes nodes in the DFD.
+One node represents the AI system, the other node represents the human oversight IT system.
+The human oversight IT system consists of all software that the human oversight personnel need to perform their duties.
+By choosing multiple processes nodes for the AI system and the human oversight IT system, we can create a top-level DFD that is valid for numerous implementations of human oversight.
+In case of threat modeling a concrete human oversight application in the future, these multiple processes nodes simply have to be specified with additional lower-level DFDs.
+Finally, the last identified node represents the human oversight personnel.
+We model this component of the human oversight application as a process to indicate the active role of humans in human oversight.
+Furthermore, the crucial role of human oversight personnel to the application is shown by the high connectivity of this node.
+We identified several data flows within our theoretical human oversight application.
+Users send inputs to the AI system, which responds by sending outputs back to the users.
+Inputs can be any data that is used to interact with the AI system.
+Similarly, outputs can be arbitrary data that represents the output of the AI system (e.g., decisions or generated data points).
+There is a data flow between users and HO personnel indicating the possibility of users to directly send feedback and reports to the HO personnel.
+Such user reports are an important oversight signal in many applications.
+Another data flow connects the AI system with the human oversight personnel.
+While the AI system send operational parameters to the human oversight personnel (this is often called monitoring in the HO literature), the oversight personnel performs interventions on the AI system [41].
+These can include commands to shut down, corrections of model parameters, or other interventions.
+The next data flow is located between human oversight personnel and human oversight IT system.
+While the personnel sends the operational parameter from the AI system to the human oversight IT system, the IT system uses these parameters to provide decision support for the oversight task.
+These will be transferred back to the human oversight personnel.
+The operational parameters could be directly send from the AI system to the human oversight IT system.
+However, This would not create a new privilege boundary crossing and, thus, can be viewed as similar to our DFD.
+We decided on the current DFD design to emphasis the importance of the human oversight personnel.
+There is a data flow between the human oversight personnel and the superior node.
+Since the superior node represents an entity of higher authority, the data flow labeling is chosen to represent this hierarchy.
+The human oversight personnel sends reports to the entity with higher authority while instructions (e.g., work instructions) are flowing the other way.
+Finally, data in form of guidelines flow from the external advisory board to the human oversight personnel.
+Five privilege boundaries can be identified in the DFD.
+First, there is a privilege boundary between users and the AI system.
+While users should have minimal rights that are necessary to use the application, the AI system needs enough rights to fulfill its task.
+Additionally, there is a privilege boundary between users and HO personnel with a similar reasoning.
+The next privilege boundary lies between the AI system and the human oversight personnel.
+Being able to intervene on the AI system requires the HO system to have higher privileges that the AI system.
+The third privilege boundary lies between the human oversight personnel and the superior node.
+Since the superior node represents management of an organization, this node has the higher privilege context.
+The fourth privilege boundary lies between the human oversight personnel and the external advisory board.
+An external advisory board has no privileges within the system.
+That means we have five different privilege levels: advisory privilege, user privilege, AI privilege, HO privilege, and superior privilege.
+3.1.2 Entry points
+Table 1: A list of all identified entry points into the human oversight application.
+Each entry point has a unique identifier, a name, and a description.
+Furthermore, the column privilege levels indicate which privilege level are required at each entry point.
+ID Entry point Description Privilege levels
+Entry.1 User login Login procedure for users to interact with the AI system.
+This could be a login page reachable from the internet using a browser.
+Other possibilities are APIs, specialized user applications, or hardware components.
+user, AI, HO, superior
+Entry.2 HO personnel login Login procedure for human oversight personnel to get access to the human oversight system.
+Possible options for this entry point are browser-based login pages, specialized applications, or hardware components.
+HO, superior
+Entry.3 Access of superiors Capabilities of entities with higher authority (e.g., management of organizations) to gain access to the human oversight system.
+This can include organizational structures and intervention capabilities.
+superior
+The next part of determining the scope of human oversight is to identify entry points into the application.
+The identification of entry points plays a crucial role in understanding possible paths that are open to a malicious actor to attack an application.
+Usually, this means to determine technical entry points like HTTP ports, login pages or functions, or search entry pages.
+Since we are not looking at a concrete implementation of human oversight, our analysis focuses on potential points of interest that need to be closely examined for real-world applications of human oversight.
+For every entry point, we also specify which privilege levels are required to use the entry point.
+Table 1 shows a summary of the entry points.
+The three identified entry points are the component used by users to access and use the AI system, the component used by human oversight personnel to access the human oversight application, and the component that allows entities with higher authority to access the human oversight application.
+3.1.3 Exit points
+Table 2: A list of identified exit points of the human oversight application.
+Each exit point has a unique identifier, a name, and a description.
+The last column indicates the corresponding entry point to each exit point, if such an entry point exists.
+ID Exit point Description Corr. entry point
+Exit.1 AI Output Communication channel that is used by the AI system to deliver outputs (e.g., predictions, decision, or generated data points) to users. Entry.1
+Exit.2 HO authentication Response functionality by the human oversight application to deliver the results of the human oversight personnel’s login procedure. Entry.2
+Exit.3 HO system report Communication channel between human oversight system and entities with higher authorities.
+This channel is used to deliver activity report and other requested information to superiors like the management of an organization.
+Entry.3
+Some attacks require an exit point to achieve their goal, i.e., a gateway that can be used to ex-filtrate data from the application.
+Two more commonly known attacks worth mentioning here are cross-site-scripting vulnerabilities and information disclosure vulnerabilities.
+In general, exit points that are not sufficiently secured can be exploited by attackers to collect sensitive information.
+From a technical point of view, exit points often have a corresponding entry point.
+All exit points are summarized in Table 2.
+We identified three exit points for the human oversight application: the AI output provided to users, the component allowing human oversight personnel to access the human oversight application, and the component used to provide work reports to entities with higher authority.
+3.1.4 Assets
+Table 3: List of identified physical assets of the human oversight application. Each asset has a unique identifier, a name, and a description.
+ID Name Description
+PA.1 HO personnel credentials The keys that human oversight personnel uses to access the human oversight system.
+PA.2 User data Depending on the concrete application, the human oversight system might store sensitive information about users, if this information is needed to determine whether an intervention has to be performed.
+To complete the first step of threat modeling human oversight, we identify assets that need to be protected.
+Assets can be both physical and abstract.
+Physical assets are for example sensitive information like medical data of patients.
+Abstract assets can be understood as important properties and states, like the reputation of an organization.
+The lists of physical and abstract assets presented in this work are not exhaustive.
+Real-world applications of human oversight are likely to have additional assets.
+Table 4: List of identified abstract assets of the human oversight application. Each asset has a unique identifier, a name, and a description.
+ID Name Description
+AA.1 Confidentiality Ensuring that sensitive information about the human oversight application, like interventions or decision protocols, cannot be accessed by unauthorized individuals.
+AA.2 Integrity Ensuring the correctness of data and functionality of the human oversight application.
+AA.3 Availability Ensuring that the human oversight application is functional at all times.
+AA.4 Epistemic access Ensuring that human oversight personnel can sufficiently understand how the AI system works.
+AA.5 Causal power Ensuring that the human oversight system can sufficiently intervene into the AI system.
+AA.6 Self-control Ensuring that the human oversight personnel has sufficient control over their own actions.
+AA.7 Fitting intentions Ensuring that the internal goals of human oversight personnel is aligned with the goals of the human oversight application.
+We identify two physical assets (Table 3).
+The first are the credentials of human oversight personnel that are used to log into the human oversight system.
+These would allow an attacker to gain access with high privileges.
+Another physical assets are sensitive user information.
+Depending on the AI system, additional information about users might be necessary to ensure that the human oversight system has sufficient knowledge about a situation to decide whether or not to intervene.
+Envision an AI system that acts as a medical decision support system.
+Judging the validity of decisions made by such a system requires detailed knowledge of patient-specific information.
+In general, such sensitive information has to be protected against unauthorized access.
+We identified seven abstract assets (Table 4).
+The first three assets that need to be protected are defined by the basic focus of cybersecurity, the CIA triad.
+This concept describes three main aspects of IT systems that have to be protected.
+The first aspect is the confidentiality of information within the system (denoted by the letter ’C’).
+Confidentiality is closely connected with the physical assets identified above but also covers appropriate management of read access within the application.
+The second aspect is the integrity of data and functionality of the IT system (denoted by the letter ’I’).
+For human oversight this can include for example ensuring correctness of AI operational parameters or correct functioning of software components.
+The last aspect covered by the CIA triad is availability (denoted by the letter ’A’).
+For example, it needs to be ensure that the human oversight personnel can access the AI system or the human oversight IT system.
+Furthermore, the obligations associated with human oversight can only be fulfilled, if such oversight can be considered effective.
+Therefore, the four requirements of effective oversight [72] have to be added to the list of abstract assets.
+The first asset is epistemic access, which means that human oversight personnel has to be able to understand the functioning of the AI system.
+The next asset is causal power, which means that the human oversight system has to be able to perform interventions on the AI system.
+The third requirement is self-control, which means that the human oversight personnel of our application can make decisions without malicious outside influence.
+Finally, the fourth requirement is fitting intentions.
+Mapping this requirement onto our human oversight application means that the goals of the human oversight personnel have to be aligned with the overall goals of human oversight.
+Based on the result of the first step of threat modeling, we now continue to identify security threats that need to be considered when implementing a human oversight application.
+3.2 Step 2 - Security threats
+We use the established threat model STRIDE (Spoofing, Tampering, Repudiation, Information disclosure, Denial of services, Elevation of privilege) to identify security threats to the human oversight application.
+STRIDE allows us to examine the application from an attacker’s perspective and to identify vulnerabilities with the attacker’s goals in mind.
+3.2.1 Spoofing
+The term spoofing summarizes threat actions that are aimed at accessing and using credentials without permission.
+The goal is to interact with a system while assuming a false identity.
+This can allow attackers to gain unauthorized access to sensitive data or functionalities.
+The obvious target for spoofing attacks in our human oversight application is the human oversight personnel and, thereby, using entry point HO personnel login (Entry.2) for exploitation.
+Human oversight personnel has to have far reaching privileges within the human oversight application, thus, gaining access through the HO personnel login can be a high priority target for attackers.
+The first set of tools readily available to attackers to pursue this target are the different methods from social engineering.
+This include water holing, baiting, phishing, vishing, smishing, and others.
+Any of the mentioned approaches can be used to steal credentials and use them for unauthorized access to the human oversight application with considerable privileges.
+If there is insufficiently secured remote access to the human oversight application, attackers can perform a man-in-the-middle (MITM) attack to get transferred information, which in this case are credentials.
+The term MITM describes attacks that rely on the attacker’s ability to secretly intercept, relay, and possibly alter communication between two parties.
+The abstract assets that are vulnerable to social engineering attacks are confidentiality (AA.1) and self-control (AA.6).
+The vulnerability of confidentiality is self-evident.
+Since the credentials themselves are considered confidential information, attacker access to credentials is a breach of confidentiality.
+The vulnerability of self-control becomes clear when looking at the goal of spoofing attacks: accessing an application while assuming a false identity.
+Attackers that use stolen credentials and are able to use these credentials to log into the human oversight application can act as human oversight personnel.
+For the system, all malicious actions taken will look like actions by human oversight personnel.
+Therefore, the human oversight personnel is no longer in control of their own actions.
+The physical asset HO personnel credentials (PA.1) is a target of these types of attacks.
+Another possible direction of spoofing attacks has emerged with the rise of more and more powerful agentic AI.
+This direction comes from the AI agents themselves.
+Research has suggested that current state-of-the-art large language models (LLMs) have the ability to do scheming [51].
+This term summarizes different misaligned behavioral patterns shown by LLMs that are meant to pursue objectives that differ from expectations of human users.
+With insufficiently secured model context protocol (MCP) servers, AI agents can try and use available tools to access credentials of human oversight personnel.
+Using these credentials, the AI agent can act within the human oversight system acting as the human oversight personnel.
+In this scenario, human oversight would be completely bypassed by the AI agent.
+Obviously, one abstract asset that is attacked in this scenario is integrity (AA.2).
+An agentic AI bypassing human oversight renders the functionality of the human oversight application useless.
+Furthermore, self-control (AA.6) is also attacked since a successful attack results in another entity being in control of the human oversight personnel’s actions.
+One might argue that also the assets confidentiality (AA.1) and HO personnel credentials (PA.1) are attacked.
+However, we would argue that misaligned behavior of AI agents do not expose confident data to the outside of the application.
+Therefore, these assets are not attacked.
+3.2.2 Tampering
+The term tampering summarizes attacks that are aimed at maliciously modifying a system.
+This can include changing persistent data, altering the system’s behavior, or messing with communication in transit.
+One target for tampering attacks is the communication from users to the AI system.
+Malicious actors can use (il)legitimately obtained user credentials to send (indirect) prompt injections to the AI system.
+Additionally, MITM attacks can be used to modify legitimate traffic coming from users with (indirect) prompt injections.
+Such attacks can cause AI systems to produce unexpected behavioral patterns.
+First asset vulnerable to this attack is integrity (AA.2) since (indirect) prompt injection cause unexpected behavior of AI systems and, thus, attack the functional integrity of the system.
+The second vulnerable asset is epistemic access (AA.4).
+Since tampering attacks alter the behavior of AI systems and can go undetected, the knowledge of the HO system on how the AI system works gets eroded.
+If the AI system has the rights to write data, the physical asset user data (PA.2) can also be vulnerable to these attacks.
+Malicious actors can try to exploit these rights and trigger the AI system to change stored user data.
+Another possible tampering attack is targeting the supply chain to poison training data or directly modify model parameters.
+Poisoning attacks can embed back doors into AI models, which can be exploited once the model is deployed.
+Again, backdooring AI models and tampering with parameters erode the HO system’s fundamental understanding of how the AI system functions.
+Thus, epistemic access (AA.4) is a target of these attacks.
+Furthermore, integrity (AA.2) is also a target, since the attacks aim to compromise the functional integrity of the AI system.
+3.2.3 Repudiation
+The term repudiation summarizes attacks that are aimed at performing prohibited operations within an IT system. These can exploit lacking capabilities to track operations within a system.
+The first attack vector that can be exploited for repudiation are adversarial attacks and (indirect) prompt injections.
+Using the communication channel between AI system and users, malicious actors can try to circumvent guardrails embedded into AI systems to trigger forbidden operations.
+In the worst case, sensitive data can be extracted exploiting the lethal trifecta [77] together with exit point AI output (Exit.1).
+If the AI system has access to log files, malicious actors can try to hide these forbidden operations.
+Therefore, (indirect) prompt injections can be used to attack both physical assets, HO personnel credentials (PA.1) and User data (PA.2).
+Furthermore, the abstract assets confidentiality (AA.1), integrity (AA.2), and epistemic access (AA.4) are vulnerable to these attacks.
+LLM’s ability to perform scheming can also open the human oversight application to repudiation attacks.
+Scheming can convert an AI system into an insider threat by triggering forbidden operations due to misaligned behavior.
+Again, these attacks can be obscured, if the AI system has the ability to alter log files.
+Epistemic access (AA.4) and integrity(AA.2) are vulnerable to this form of attack, since scheming prevents human oversight personnel to fully understand the functioning of the AI system and the functionality of the AI system is compromised.
+Depending on how extensive the misaligned AI system can operate within the application, the asset causal power can also be vulnerable.
+An AI system with sufficient rights can potentially limit the intervention capabilities of the human oversight personnel.
+Another set of attack vectors for repudiation attacks are directly targeted at the human oversight personnel.
+Bribery and coercion can be used by malicious actors to force human oversight personnel to perform prohibited operations.
+Again, being able to alter the log files allows human oversight personnel to hide these forbidden operations.
+The abstract assets confidentiality (AA.1) and integrity (AA.2) are vulnerable to these types of attacks, since both bribery and coercion can be used to extract sensitive information from the system and maliciously alter persistent data and functionalities.
+Furthermore, coercion attacks self-control (AA.6) due to the outside force that determines the actions for the human oversight personnel.
+Bribery attacks fitting intentions (AA.7), since the original goal of the human oversight personnel gets overridden with monetary or otherwise self-serving interests.
+3.2.4 Information disclosure
+The term information disclosure summarizes attacks that aim at reading files without permission or reading data in transit.
+Since all attacks that fall into this category are explicitly targeting the disclosure of sensitive information, the asset that is mainly targeted by these attacks is confidentiality (AA.1).
+Furthermore, the two identified physical assets HO personnel credentials (PA.1) and user data (PA.2) are sensitive information and, thus, can also be targeted by information disclosure attacks.
+The two data flows that will be the main target of information disclosure attacks are between AI and human oversight personnel, and human oversight personnel and IT system.
+Both data flows contain valuable information about the inner workings of the human oversight applications and, hence, might be interesting for attackers.
+The data flow between users and AI system could also be attacked.
+However, this data flow contains mainly information about the inner workings of the AI system.
+This data flow plays only a secondary role for the threat modeling of human oversight.
+A standard attack vector for data in transit are MITM attacks.
+3.2.5 Denial of services
+The term denial of services covers attacks attempting to deny access to valid users, such as by making a web server temporarily unavailable or unusable.
+One of the most well-known attack patterns that fall into this category is flooding a server with more requests than can be reasonably processed.
+This leads to the server becoming unresponsive.
+These attack patterns are also called (distributed) denial of services ((D)DOS) attacks.
+In general, threat actions that fall under the term denial of services are targeting the abstract asset availability (AA.3).
+The first obvious target for denial of services attacks is the human oversight IT system.
+If the network that runs the IT system is accessible from the internet, malicious actors can deploy (D)DOS attacks to flood the network with requests.
+Another possibility is the exploitation of software vulnerabilities to cause a shutdown of the human oversight IT system.
+Since the human oversight IT system provides help to human oversight personnel to understand the AI system’s behavior and decision support, a denial of service attack on the IT system additionally targets the abstract asset epistemic access (AA.4).
+In a distributed setting, where the HO system is not co-located with the AI system, malicious actors can specifically target the network of the human oversight personnel.
+This would prevent human oversight personnel from intervening into the AI system and, thus, additionally targets the abstract asset causal power (AA.5).
+3.2.6 Elevation of privilege
+The term elevation of privilege summarizes attacks that aim at escalating the own privileges within a system.
+An elevation of privileges attack uses vulnerabilities and design/configuration flaws after an initial network breach to get access to data and functionalities that are normally inaccessible for the initial privilege group.
+Due to our extended definition of privilege, an elevation of privilege attack can also lead to intervention rights that normally would not be accessible.
+An attack vector that can be exploited by malicious actors to attack the AI system within the human oversight application is jailbreaking.
+Here, different techniques like poisoning attacks and adversarial attack can be used to circumvent guardrails that prevent AI systems from misaligned behavior.
+Together with the exploitation of software and design flaws this can lead to the AI system having similar privileges than the human oversight components.
+In this case, interventions issued by the human oversight personnel could be overwritten by the AI system.
+Therefore, these attacks target the abstract asset causal power (AA.5)
+3.3 Step 3 - Countermeasures and hardening strategies
+The STRIDE analysis revealed several vulnerabilities within the human oversight application.
+The next step in our threat modeling highlights a few countermeasure and mitigation strategies that should be adhered to when implementing human oversight.
+The list is not exhaustive since our theoretical threat modeling cannot determine all threats that an actual implementation might face.
+For example, there might be mitigation strategies that depend on specific hard- and software components installed into the human oversight application.
+Therefore, the provided list can be understood as a starting point and minimal set.
+3.3.1 Intrusion Detection System
+Several threat actions identified in section 3.2 require an initial breach of the application environment.
+Therefore, the application’s environment, including all computer and network components, should be closely monitored using intrusion detection systems (IDS).
+These systems typically combine multiple components and methods to perform this monitoring.
 The core technical components of an IDS include network or host sensors (depending on the IDS type), a database, a management module, and an analysis module.
@@
 A well-designed and maintained IDS can significantly enhance the cybersecurity of IT systems [8, 36].
-Intrusion detection systems may counter multiple attack vectors, including (D)DoS attacks, MitM attacks, malware, and insider threats.
-Hardening strategy: Organizations should implement intrusion detection systems to protect their networks broadly and specifically to safeguard human oversight operations and the tools used by oversight personnel from malicious actors.
-4.3 Encryption
-Encrypted communication is the standard for data traffic on the internet.
-However, non-encrypted or weakly encrypted data traffic is still commonly used, particularly for communications that do not involve direct interaction between a user and a website.
-Such vulnerable traffic enables malicious actors to carry out Man-in-the-Middle (MitM) attacks relatively easily, often without requiring technical expertise or expensive software and hardware.
+3.3.2 Encryption
+Another mitigation strategy that protects sensitive information is encryption.
+Non-encrypted data traffic enables malicious actors to carry out for example Man-in-the-Middle (MitM) attacks with relative ease, often without requiring technical expertise or expensive software and hardware.
+This makes an application vulnerable to spoofing, tampering, and information disclosure threats.
 To ensure adequate security, modern cryptographic methods with strong security guarantees must be employed to encrypt data traffic [38, 20].
-Hardening strategy: All communication between human oversight personnel and the AI system must be encrypted.
-When personnel and the AI system are not co-located, using a virtual private network to protect data traffic is recommended.
-4.4 Transparency
-One major challenge in securing AI systems is their complexity.
-Enhancing transparency of both the system and its communication with human oversight personnel can improve cybersecurity by facilitating the detection of attacks.
+Furthermore, encrypted data storage should be considered throughout the human oversight application.
+3.3.3 Network Management
+As indicated in section 3.2.5, denial of service threats overwhelm a network by generating excessive traffic.
+Defending against these forms of attacks requires a multi-step approach [56, 11].
+The first step is to deploy monitoring systems that analyze network traffic and trigger alerts if a potential (D)DoS attack is detected.
+The second step involves methods to distinguish between benevolent traffic (typically generated by humans) and malicious traffic (often produced by bots).
+Finally, filtering techniques can mitigate (D)DoS attacks by restricting or limiting what users – humans or bots – can do within the network.
+3.3.4 Transparency
+Enhancing transparency can improve cybersecurity by facilitating the detection of attacks.
 While transparency itself is not typically considered a cybersecurity measure, it provides access to key security-relevant information, such as implemented components, training data, system functionality, development processes, and dependencies.
@@
 Additionally, transparency supports auditing and compliance attestations by making security-relevant information easily accessible and enables model tracking.
-Overall, ensuring transparency can also help human oversight personnel’s response time, as decision-relevant information becomes more readily available.
-An appropriate level of transparency can help to counter all of the attack vectors mentioned in this work.
-Hardening strategy: Promote transparency to enhance attack detection by making security-relevant information accessible and making potential manipulations of the system easier to identify.
-4.5 Governance Principles and Audit Frameworks
-Governance principles and audit frameworks are essential for effectively managing security risks in highly automated systems.
-Governance principles provide the fundamental guidelines, rules, and processes that define how an organization makes decisions and ensures accountability.
-They establish the framework for achieving strategic goals while minimizing ethical, legal, and operational risks.
-In the context of AI, these principles specify who is responsible for risk assessment, how transparency is ensured, and which standards apply to data security.
-A prominent example of implementing these principles is the AAA Framework, which stands for Prospective Assessments, Audit Trail, and Enforceable Compliance.
-These three pillars offer a comprehensive approach to strengthening system integrity and human oversight.
-Prospective Assessments, conducted before system implementation, identify and mitigate potential risks early, incorporating the perspectives of affected individuals and groups [66].
-A robust Audit Trail, supported by high-quality data and system transparency, is essential for analyzing errors and assigning responsibility.
-Enforceable Compliance ensures that legal requirements are actionable, with independent audits verifying compliance and enhancing confidence in system integrity.
-Another well-established framework is the information security management system (ISMS) specified by ISO/IEC 27001 [92].
-An ISMS provides the technical and organizational foundation to ensure integrity and reliability of information such as data, commands, and system statuses.
-Governance principles and audit frameworks are a countermeasure against vulnerability exploitations and insider threats.
-Hardening strategy: Organizations need to establish governance principles and audit frameworks after established standards like the AAA framework and ISO/IEC 27001.
-These principles and frameworks have to include all components of human oversight that are implemented within an organization.
-4.6 Red Teaming
-The term red team refers to an independent group that investigates an organization’s IT systems to identify vulnerabilities and security flaws [93].
-The red team simulates the behavior of malicious actors, employing all methods that real attackers might use.
-Its goal is to uncover as many exploitable security issues within the tested system as possible.
-Red teaming can help counter several different attack vectors, including poisoning attacks, adversarial attacks, explainability attacks, and vulnerability exploitation.
-Hardening strategy: Organizations should conduct regular red teaming exercises to detect and address security flaws within their human oversight processes.
-4.7 Training of Human Oversight Personnel
+Transparency can also help to secure the supply chain of AI, for example with an SBOM for AI [21].
+This could help as a countermeasure against AI specific attacks, like poisoning or adversarial attacks.
+3.3.5 Training of Human Oversight Personnel
 Training aims to strengthen the resilience of human oversight personnel against attacks [64].
@@
 Practical exercises, such as simulated phishing campaigns or role-play scenarios, reinforce recognition and proper reporting behaviors.
-Training can also address resilience against coercion, preparing personnel to respond to threats or undue pressure.
-This includes protocols for escalating coercion attempts, access to secure whistleblowing channels, and preparation to maintain effective human oversight behavior under stress.
-Anti-bribery training can teach personnel to recognize undue influence.
+It can also address resilience against coercion, preparing personnel to respond to threats or undue pressure.
+This includes protocols for escalating coercion attempts, access to secure whistleblowing channels, and preparation to maintain effective human oversight under stress.
+Anti-bribery training can teach recognizing undue influence.
 Clear organizational policies defining conflicts of interest and mandatory disclosure mechanisms further reduce the risk of misaligned intentions.
 Training of human oversight personal counters social engineering, coercion, and bribery.
-Hardening strategy: Organizations should implement recurring, context-specific training programs for human oversight personnel.
-By institutionalizing such training and embedding it in the organization’s security culture, human oversight personnel can develop the awareness, self-efficacy, and procedural support needed to counteract manipulation attempts, thereby contributing to secure human oversight.
+3.3.6 Red Teaming
+The term red team refers to an independent group that investigates an organization’s IT systems to identify vulnerabilities and security flaws [49].
+The red team simulates the behavior of malicious actors, employing all methods that real attackers might use.
+Its goal is to uncover as many exploitable security issues within the tested system as possible.
+Continuous red teaming can help to stay on top of emerging security flaws after the initial threat modeling of an application.
+Software systems are dynamic even in deployment and, thus, require constant monitoring.
 4 Discussion
@@
 In this paper, we highlight the importance of complementing this focus with a security-related perspective on human oversight.
-This perspective is critical because various attack vectors can undermine human oversight effectiveness, a risk that will grow as human oversight becomes more widely implemented as part of the safety, security, and accountability architecture of AI operations.
-Drawing on cybersecurity research [33, 44], we have introduced various attack vectors targeting human oversight, along with corresponding hardening strategies, to support the development of secure human oversight of AI.
+This perspective is critical because digital systems are increasingly pressured by cyber criminals and weaknesses in single components can compromise whole IT infrastructures, a risk that will also apply to human oversight as it becomes more widely implemented as part of the safety, security, and accountability architecture of AI operations.
+Drawing on cybersecurity research [65, 15], we have introduced threat modeling for human oversight.
+We designed a DFD for the human oversight application, identified assets that need to be protected, and explored vulnerabilities of human oversight using STRIDE.
+Furthermore, we introduced a minimal set of mitigation strategies to secure human oversight.
 Understanding and designing the technical factors, selecting and training human oversight personnel, and shaping working conditions to foster effective human oversight remains an ongoing challenge for research, practice, and AI governance [72, 25, 44].
 Our analyses indicate that this challenge will intensify as human oversight becomes standard practice in high-risk AI use cases.
-Malicious actors targeting the attack surface of human oversight can compromise the AI system, the communication between humans and AI, or the oversight personnel themselves, aiming to undermine epistemic access, causal power, self-control, and fitting intentions as the core requirements for effective oversight [1].
+By analyzing an abstracted application-level representation of human oversight using structured threat modeling, we unveiled vulnerabilities that can be exploited by malicious actors to attack key assets, including the requirements of effective human oversight [72].
 Consequently, such attacks may increase risks in AI operations by reducing the effectiveness of oversight, disabling oversight entirely, or manipulating oversight personnel’s intentions to exacerbate risks.
-The attack vector list we present is not exhaustive; future research should explore novel attack vectors specific to human oversight.
-A potential future scenario involves the deployment of malicious AI agents designed to deceive oversight [70, 69].
-Once an AI system models the oversight mechanism, it may attempt to circumvent the restrictions imposed by human oversight, further expanding the attack surface and associated risks.
-Given these risks, it is essential to consider how to enhance the security of human oversight.
-To this end, we presented various hardening strategies drawn from the broader cybersecurity literature [52, 53, 54].
-These strategies must be actively applied and adapted to the context of human oversight of AI.
-Our paper contextualizes these strategies for human oversight, representing a crucial first step toward securing oversight architectures against malicious attacks.
-Future research needs to evaluate the effectiveness of these hardening strategies and could develop additional strategies tailored specifically to human oversight.
+We were able to identify a list of security risk using the STRIDE model.
+However, this list can only serve as a starting point for real applications of human oversight.
+In practice, our baseline has to be extended with the concrete software and hardware components that are integrated into the human oversight application.
+These components can open an application to new attack vectors.
+Furthermore, we included threats posed by misaligned AI agents into our investigation.
+With growing knowledge about AI agents in practice, new ways possibly emerge how AI agents can deceive oversight [51, 7].
+While we were using STRIDE to analyze our theoretical human oversight application, additional options are available for the secure implementation of human oversight.
+Once human oversight is operational, the analysis with STRIDE should be supplemented with operational threat modeling using frameworks like cyber kill chain [79], MITRE ATT&CK [4], or unified kill chain [59].
+This allows for an even more comprehensive and specific analysis of the operational aspects of a human oversight application.
+Another endeavor that should be considered by practitioners when performing real-world threat modeling is the quantitative analysis of threats.
+This means determining for each threat the ease of exploitation and the damage potential of an exploitation.
+Such analysis can help to prioritize hardening strategies.
+Identifying risk is only one side of the cybersecurity coin.
+Appropriate hardening strategies have to be put into place to secure applications.
+We introduce several hardening strategies that can be used to mitigate the identified vulnerabilities.
+These strategies are based on the cybersecurity literature [62, 61, 9].
+However, our provided list has to be understood as a minimal set of measures.
+Future research needs to identify additional measures and evaluate the effectiveness of common hardening strategies for human oversight.
 Beyond technical hardening, future work could also explore the application of amplified oversight.
@@
 In future, such approaches to support human oversight may even exceed human capabilities and may become a further viable target for malicious actors to attack.
-While this paper focuses on malicious actors targeting human oversight, not all risks emerge from malicious actors.
-Additional risks can for example arise from counterproductive behavior within organizations [97].
-Counterproductive behavior refers to intentional or unintentional actions by employees that go against organizational goal, such as neglecting oversight duties, resisting established procedures, or even deliberately undermining AI operations, which can erode the effectiveness of human oversight.
-Such behavior may stem from factors like dissatisfaction with working conditions, misaligned incentives, organizational conflict, or lack of recognition [97].
-Such counterproductive behavior inside organizations may need closer examination in the future to ensure secure human oversight.
-Such behavior can undermine the core requirements of effective oversight and may promote active damage to the organization or passive under-performance, thereby introducing risks to human oversight of AI.
-Although these behaviors are not attacks by malicious actors, they may make attacks by external actors more effective, for example, if an attacker targets a dissatisfied employee.
-Future works could shed light on those risk factors and suitable hardening strategies such as organizational structures, incentive systems, and thoughtful work design [29].
-Well-designed security concepts, policies and work conditions can all together mitigate security risks and prevent counterproductive behavior by ensuring that oversight personnel’s goals remain aligned with organizational objectives.
 5 Conclusion
@@
 Adding to this complexity, it will become increasingly important to design for secure human oversight as oversight becomes an integral component of the safety, security, and accountability infrastructure of AI operations.
-We hope this paper serves as a starting point for research that analyzes the attack surface of human oversight and develops effective hardening strategies to prevent malicious actors from exploiting oversight as a gateway to compromise AI safety.
+We hope this paper serves as a starting point for research that analyzes the cybersecurity risk of human oversight.
+Additionally, we hope our paper serves as a guideline for researchers and technicians that are task to implement human oversight in a real-world system.
 Acknowledgments